### v1

# ChartMimic: Evaluating LMM’s Cross-Modal Reasoning Capability via Chart-to-Code Generation

Abstract
We introduce a new benchmark, ChartMimic, aimed at assessing the visually-grounded code generation capabilities of large multimodal models (LMMs). ChartMimic utilizes information-intensive visual charts and textual instructions as inputs, requiring LMMs to generate the corresponding code for chart rendering. ChartMimic includes $1,000$ human-curated (figure, instruction, code) triplets, which represent the authentic chart use cases found in scientific papers across various domains (e.g., Physics, Computer Science, Economics, etc). These charts span $18$ regular types and $4$ advanced types, diversifying into $191$ subcategories. Furthermore, we propose multi-level evaluation metrics to provide an automatic and thorough assessment of the output code and the rendered charts. Unlike existing code generation benchmarks, ChartMimic places emphasis on evaluating LMMs’ capacity to harmonize a blend of cognitive capabilities, encompassing visual understanding, code generation, and cross-modal reasoning. The evaluation of $3$ proprietary models and $11$ open-weight models highlights the substantial challenges posed by ChartMimic. Even the advanced GPT-4V, Claude-3-opus only achieve an average score of $73.2$ and $53.7$, respectively, indicating significant room for improvement. We anticipate that ChartMimic will inspire the development of LMMs, advancing the pursuit of artificial general intelligence.
1 Introduction
Code generation [1] is a rather demanding task that requires advanced abstract thinking and logical reasoning, reflecting the unique intelligence of human beings. Recently, progresses in artificial general intelligence (AGI) have demonstrated the potential of large foundation models [2, 3] in solving these tasks that were once the exclusive domain of human abilities [4, 5]. However, existing code generation benchmarks [6, 7, 8, 9] solely use text as input, while humans receive information from multiple modalities when coding [10, 11], such real-life scenario has not been fully explored.
Taking a common scene in Fig. 1, researchers often need to write code for data visualization and may already have preferred chart templates at hand. However, they usually lack either the source code or the expertise to reproduce these chart templates. As a result, they turn to large multimodal models (LMMs) as assistants to aid in code generation. In this scenario, coding for scientific charts entails code generation grounded on visual understanding (i.e., chart-to-code generation), which necessitates LMMs to integrate a variety of advanced cognitive capabilities, including visual understanding, code generation, and cross-modal reasoning. Therefore, evaluating the performance of LMMs on this real-world task also enables researchers to pinpoint potential areas for improving models’ capabilities.
To this end, we present ChartMimic (Fig. 2), a multimodal code generation benchmark. ChartMimic is characterized by its (1) information-intensive visual inputs, (2) diverse chart types and (3) multi-level evaluation metrics. Specifically, compared to natural images, scientific charts convey nuanced semantic meanings through intricate visual logic, thereby exhibiting higher information density. Based on this, we define two tasks, Direct Mimic and Customized Mimic (Sec. 2.1), utilizing charts and textual instructions as inputs. These tasks challenge LMMs to generate the corresponding code for a given chart or to incorporate new data specified in the instructions, respectively. Through the collection of academic documents and scientific papers, we identify $22$ commonly used chart types and $191$ subcategories. Subsequently, we manually annotate a total of $1,000$ (figure, instruction, code) triplets for these types (Sec. 2.2). Furthermore, we establish automatic evaluation metrics from both high-level and low-level perspectives to thoroughly assess the performance of LMMs (Sec. 2.4).
We conduct examination of $14$ LMMs on ChartMimic (Sec. 3.2), including $3$ proprietary models and $11$ open-weight models. We observe that while several open-weight models can match the performance of proprietary models such as GPT-4V on public leaderboards [12], a significant performance gap still persists on ChartMimic. Specifically, the best open-weight model, Phi-3-Vision, significantly lags behind GPT-4V, with only half the performance on two tasks, indicating substantial room for improvement. Additionally, correlation analysis (Sec. 4.3) demonstrates the high correlation between our automatic metrics and human judgment. Further error analysis (Sec. 4.4) reveals that hallucination notably hinders LMMs’ performance on ChartMimic, involving the insertion of non-existent text in the ground-truth figures and confusion of similar chart types.
Figure 1: The real-world example. LMMs assist scientists and researchers in understanding, interpreting and creating charts during the reading and writing of academic papers. These models serve as assistants that enhance the comprehension and presentation of data in scholarly communications.
We envision ChartMimic as a suite of benchmarks designed to guide researchers in understanding the capabilities of their LMMs. By providing a comprehensive evaluation framework, ChartMimic aims to facilitate the growth of foundation models for the community, offering insights into LMM’s various aspects, such as visual understanding, code generation, and cross-modal reasoning.
2 The ChartMimic Benchmark
In this section, we first introduce the definition of two tasks involved in ChartMimic, and then delineate the data curation process. Subsequently, we conduct a quantitative analysis to demonstrate the quality and diversity of ChartMimic and compare it with existing related benchmarks.
2.1 Task Definition
The capability of LMMs to generate code that can accurately render a specific chart demonstrates their proficiency in visual understanding, code generation and cross-modal reasoning. Specifically, given the chart $X$ and the instruction $I$, the LMM $f$ is expected to output the code $C$ that satisfies the requirements outlined in the instruction:
$C=f(X,I).$ (1)
As shown in Fig. 2, based on the information provided in the instructions, we propose two tasks:
Figure 2: The pipeline of ChartMimic. We provide $1,000$ human-curated (figure, instruction, code) triplets. We use ChartMimic to evaluate LMMs’ proficiency in the multimodal chart-to-code generation, resulting in both high-level and low-level evaluation results.
Direct Mimic. The LMM is requested to directly generate code that can reproduce the provided chart. This task evaluates the LMM’s ability to comprehensively understand and reason about the chart to generate executable code.
Customized Mimic. With additional customized data provided in the instruction, this task requires the LMM to generate code that can render a new chart encoding the customized data while maintaining the aesthetic and design of the provided chart. This task requires the LMM to understand both the chart and customized data to integrate them effectively.
After obtaining the code generated by LMMs, we execute it to render the corresponding chart and subsequently compare its similarity with the ground-truth chart. Illustrative examples of the two tasks are shown in Fig. 2. To accommodate the defined tasks above, we propose ChartMimic, a benchmark designed to evaluate the comprehension of charts and their conversion into executable code.
2.2 Data Curation Process
ChartMimic distinguishes itself through $4$ fundamental considerations: (1) diversity of chart types, (2) balance of charts complexity, (3) reduction of data leakage, (4) integration of authentic user requirements. We keep on these principles to complete the data curation for ChartMimic through a four-step pipeline. We provide an overview here and more details in Appendix A.
General Filtering. We scrape figures from source files of publications on arXiv [^1] https://arxiv.org/ that hold a CC BY 4.0 license and have a publication date after February 2024. We extract figures in PDF format, yielding approximately $174,100$ figures across various domains (e.g., Physics, Computer Science, Economics, etc). We then filter the figures based on the criteria of how designers select inspiring visualization examples, including clarity and visual appeals [13, 14], color schemes [13], and uniqueness of the chart within its category, resulting in a refined collection of about $15,800$ figures.
Diversity and Information Density Filtering. We first initialize a data pool, creating a taxonomy of chart types and subcategories and comparing each chart against the most similar one within the corresponding category in the data pool. Should significant differences in visual complexity or information density [13] be found, the chart is added to the pool. Otherwise, we retain the one that better fulfills the criteria for diversity and effective information communication. Ultimately, this approach yields a curated collection of $279$ charts from the generally filtered $15,800$ ones.
Manual Selection and Filling. In addition to sourcing from arXiv, we also collect charts from various platforms, including the Matplotlib gallery, Stack Overflow, and plotting-related forums on Twitter and Reddit. These charts are selected for their unique styles not represented in our arXiv curated $279$ charts. To mitigate the risk of data leakage, we rigorously process the data and color styling of these charts, replacing existing color schemes with those unseen in our data pool while maintaining their aesthetic appeal. Consequently, we obtain $500$ prototype charts for ChartMimic.
Code and Instruction Writing. We propose to manually write codes and instructions for ChartMimic based on the collected 500 prototype charts. Initially, the annotators are tasked with meticulously reproducing the 500 prototype charts using Python code, resulting in a set of 500 (figure, code, instruction) triplets for the Direct Mimic task. Subsequently, to simulate the scenario of Customized Mimic, annotators are instructed to modify the chart data in the Direct Mimic task by integrating new data from various domains. They are then required to modify the corresponding codes and instructions, leading to the 500 (figure, code, instruction) triplets for the Customized Mimic task. Consequently, we obtain the ChartMimic benchmark, comprising $1,000$ high quality test examples.
Table 1: Statistics of ChartMimic. Each one in Combination, HR, Muiltidiff, and PIP is an independent subcategory. We measure code length in terms of tokens, utilizing the Llama3 tokenizer [15]. In the level count, “A/B/C” denotes the number of chart for “easy/medium/hard” level, respectively.
Type Bar Line ErrorBar Heatmap Box Scatters Hist Radar 3D Pie ErrorPoint Violin
Count 100 80 30 30 25 25 20 20 15 15 10 10
Subcategories 16 8 12 4 6 4 3 6 5 8 5 3
Code Length (AVG.) 689.3 805.7 651.9 664.6 659.9 617.2 507.3 772.2 668.7 385.1 523.7 912.0
Code Length (STD.) 289.0 272.2 146.1 303.0 248.9 287.6 161.2 136.4 168.9 123.1 228.9 298.7
Level Count 53 / 40 / 7 74 / 6 / 0 17 / 13 / 0 0 / 19 / 11 15 / 10 / 0 20 / 5 / 0 13 / 7 / 0 13 / 7 / 0 2 / 8 / 5 10 / 5 / 0 7 / 3 / 0 4 / 6 / 0
Type Area Contour Density Graph Quiver Treemap Combination HR Muiltidiff PIP Total
Count 5 5 5 5 5 5 30 25 25 10 500
Subcategories 2 3 4 4 4 4 30 25 25 10 101+(90)
Code Length (AVG.) 832.8 352.8 431.8 310.8 845.0 302.2 641.9 689.9 788.1 958.9 682.9
Code Length (STD.) 177.5 136.7 71.9 61.7 607.0 34.7 159.0 301.5 331.0 285.9 285.1
Level Count 4 / 1 / 0 0 / 0 / 5 3 / 2 / 0 5 / 0 / 0 0 / 2 / 3 5 / 0 / 0 3 / 19 / 8 1 / 4 / 20 0 / 12 / 13 0 / 0 / 10 249 / 169 / 82
2.3 Data Statistics
Notably, the (figure,code,instruction) triplets for both Direct and Customized Mimic tasks share the same figure. Therefore, we detail the data statistics for the $500$ triplets in the Direct Mimic task. As depicted on the left side of Fig. 2, ChartMimic encompasses a total of $22$ categories, with $18$ types of regular charts and $4$ types of advanced charts. For the $18$ regular chart types, we identify $101$ subcategories, with the definitions and examples of each subcategory provided in Appendix B. The advanced chart types such as Plot-in-Plot (PIP), Multidiff, and Combination are distinct forms of amalgamating multiple chart sets into a singular chart. Given the variety of their internal combination elements, each of these can be treated as a unique subcategory. Meanwhile, the Hard-to-Recognize (HR) category encapsulates unclassifiable charts, with each chart being considered a category unto itself. When we factor in the additional $90$ subcategories represented by these advanced chart types, ChartMimic encompasses a total of $191$ subcategories. This extensive diversity underscores the comprehensive nature of our benchmark. Additionally, we employ Llama3 [15] tokenizer to measure the code length. The results are shown in Table 1, which shows an average code token length of $635.98$ with a standard deviation of $253.06$. Moreover, we conduct a manual complexity assessment for each chart, classifying them into three levels: easy, medium, and hard. The distribution of charts according to complexity levels is as follows: $249$ for easy, $169$ for medium, and $82$ for hard.
2.4 Evaluation Metrics
For tasks within ChartMimic, an appropriate evaluation necessitates comparing the visual similarity between the generated and ground-truth figures. To achieve automatic evaluation, we propose multi-level metrics (i.e., low-level and high-level), which assess the similarity at different granularities. We compute the average score between the low-level and high-level scores, which range from 0 to 100, as the overall score. The illustration of the multi-level metrics is depicted in Fig. 2.
GPT-4V Score. Following the successful utilization of large foundation models like GPT-4 for automated evaluation in both natural language processing [16, 17, 18] and computer vision [19, 20], we adopt the GPT-4V score as our high-level metric. Specifically, we input both the ground-truth chart and the generated chart to GPT-4V, and instruct GPT-4V to output a high-level similarity score between 00 and $100$. The detailed prompt template can be found in Appendix C.
Table 2: A comparison of our proposed ChartMimic to other benchmarks. “I” and “NL” indicate “Image” and “Natural Language” respectively.
Benchmarks Source # of Chart Types # of Test Instances Input Format Output Format Evaluation Metric
Chart Understanding Benchmarks
ChartQA [21] Human Curated 3 10K I+NL NL Accuracy
Chart-to-Text [22] Crawl 6 44K I+NL NL Match-based
ChartSumm [23] Human Curated 3 84K I+NL NL Match-based
Code Generation Benchmarks
HumanEval [6] Human Curated - 164 Code Code Pass Rate
MBPP [7] Human Curated - 500 NL+Code Code Pass Rate
MMCode [24] Crawl - 263 I+NL Code Pass Rate
MatPlotBench [20] Human Curated 13 100 NL Code GPT-4V Score
Design2Code [25] Crawl HTML 484 I+NL Code Multi-Level
ChartMimic (Ours) Human Curated 22 1K I+NL Code Multi-Level
In addition to high-level similarity, evaluating low-level elements between generated and ground-truth figures can help better analyze the behaviors of LMMs. Therefore, we propose evaluating four key low-level elements: text, layout, type and color. Extracting them from figures is a challenging task, as existing extraction models often fall short in terms of accuracy [26]. Considering that figures are rendered based on the code, we design a code tracer to monitor the execution process of the ground-truth code and generated code. The code tracer records the text, layout, type and color information. We calculate the F1 score of these elements as their corresponding score. And we average text, layout, type and color scores to get the low-level score. The methodology for obtaining these elements is briefly introduced as follows, with detailed descriptions provided in Appendix C.
Text Score. During the code execution process, for the function responsible for adding text elements to the rendered figures, the code tracer monitors it and records each text element parameter.
Layout Score. Layout refers to the arrangement of subplots in the figure. At the end of the code execution, the code tracer traverses all subplots in the figure and obtains their layout information.
Type Score. For each plot function, which serves the purpose of adding a specific chart type instance to the figure, the code tracer monitors its invocation status and records every invoked plot function.
Color Score. For each plot function, it will return the chart type instance at the end of the function invocation. The code tracer accesses the color attributes of these chart type instances.
2.5 Comparisons with Existing Benchmarks
To further distinguish the difference between ChartMimic and other existing ones, we elaborate the benchmark details in Table 2. From the chart understanding perspective, the prior benchmarks [21, 22, 23] are heavily focused on simple questions about the data in the charts and lack of both difficulty and diversity in chart types. In the code generation aspect, HumanEval [6], MBPP [7] and MatPlotBench [20] only consider tasks with text inputs, which may not meet requirements in the era of LMMs. Recently, MMCode [24] attempted to create a benchmark for multimodal code generation, but the vision inputs for their task are still overly simple and only have a single pass rate evaluation metric. Design2Code [25] is the most similar one to ours, which focuses on automated webpage generation. Although they use multi-level evaluation metrics like ours, their test HTMLs all come from the C4 dataset [27], which may pose a risk of data leakage. Our ChartMimic benchmark gives complex scientific charts as inputs, demanding capabilities on grounding visual understanding into code generation, and provides multi-level evaluation metrics.
3 Experiment
3.1 Baseline Setup
We benchmark $14$ widely utilized proprietary and open-source models currently available in the field. For proprietary models, we consider $3$ representative models: GPT-4V [2], Claude-3-opus [28] and GeminiProVision [3]. For the open-weight models, we choose $11$ competitive models. We list them in increasing order of total parameter size from 4.2B to 34.8B, which are: Phi-3-Vision [29], LLaVA-Next-Vicuna-7B [30], DeepSeek-VL-7B [31], LLaVA-Next-Mistral-7B [30], IDEFICS2-8B [32], MiniCPM-Llama3-V 2.5 [33], Qwen-VL-Chat [34], LLaVA-Next-Vicuna-13B [30], Cogvlm2-llama3-chat-19B [35], InternVL-Chat-V1.5 [36], LLaVA-Next-Yi-34B [30]. We start with the direct prompting method, which provides the reference chart with direct instructions. The specific instructions and models’ decoding setups can be found in Appendix D.
3.2 Main Results
Table 3: The ChartMimic leaderboard with Direct Mimic task. The best scores are in bold. We also include the code execution success rate (Exec. Rate) and model size (Params).
Model Params Exec. Rate Low-Level High-Level Overall
Text Layout Type Color Avg. GPT-4V
Proprietary
GeminiProVision - 64.2 36.9 58.6 45.1 39.6 45.0 38.1 41.6
Claude-3-opus - 86.4 61.8 79.0 45.3 37.9 56.0 45.4 50.7
GPT-4V - 91.4 74.3 87.1 73.8 62.1 74.3 68.4 71.4
Open-Weight
Qwen-VL-Chat 9.6B 4.0 0.5 2.8 1.1 0.8 1.3 0.9 1.1
LLaVA-Next-Vicuna-7B 7.1B 6.2 0.8 2.1 1.4 1.6 1.5 1.6 1.6
LLaVA-Next-Vicuna-13B 13.4B 11.6 1.9 5.7 4.0 4.3 4.0 2.8 3.4
IDEFICS2-8B 7.6B 42.4 5.7 28.4 11.9 10.9 14.2 8.8 11.5
LLaVA-Next-Yi-34B 34.8B 45.6 11.4 26.7 18.2 16.6 18.2 12.7 15.4
LLaVA-Next-Mistral-7B 7.6B 57.6 11.3 31.5 22.3 21.9 21.8 14.8 18.3
DeepSeek-VL-7B 7.3B 39.0 13.8 26.4 23.9 19.0 20.8 16.8 18.8
Cogvlm2-llama3-chat-19B 19.2B 51.4 20.6 31.1 21.7 20.0 23.3 17.4 20.4
InternVL-Chat-V1.5 26.0B 60.0 26.9 42.1 26.1 24.9 30.0 23.4 26.7
MiniCPM-Llama3-V2.5 8.4B 58.2 22.5 40.1 31.4 26.6 30.2 24.2 27.2
Phi-3-Vision 4.2B 66.0 36.4 45.0 39.0 34.3 38.7 25.1 31.9
We present the main results of 14 LMMs on ChartMimic. Table 3 and Table 4 show the results on the Direct Mimic and Customized Mimic task, respectively. The key findings are as follows:
GPT-4V performs best among proprietary models, while Phi-3-Vision excels among open-weight models. In proprietary models, GPT-4V achieves an overall score of $71.4$ in Direct Mimic and $72.33$ in Customized Mimic, surpassing the second-place model Claude-3-opus by $21.3$ and $21.1$, respectively. Within open-weight models, Phi-3-Vision, despite its 4.2B parameters, reaches the overall score of $31.9$ in Direct Mimic and $40.18$ in Customized Mimic, which outperforms LLaVA-Next-Yi-34B ($15.4$ in Direct Mimic and $20.2$ in Customized Mimic). This indicates that even models with fewer parameters can achieve decent performance through a refined training process.
A performance gap persists between open-weight LMMs and proprietary ones. Although several open-weight LMMs can exhibit performance comparable to GPT-4V across various benchmarks [12], even the best-performing Phi-3-Vision falls short of achieving half the performance of GPT-4V on both Direct Mimic and Customized Mimic. This apparent performance disparity demonstrates the challenging nature of our ChartMimic benchmark for current open-weight LMMs. Moreover, open-weight LMMs even exhibit notable deficiencies in generating executable code, with a majority of them showing an execution rate below $60\%$. These findings emphasize that there is still a considerable scope for the open-source community to enhance LMMs’ capabilities in terms of complex visual understanding, code generation and cross-modal reasoning.
When given additional user-customized data, LMMs exhibit performance improvements. In the Customized Mimic task, most LMMs get performance improvement compared to the Direct Mimic task, particularly in terms of Text Score. This enhancement can be attributed to the provision of user-customized data, which alleviates the burden on LMMs to recognize textual and data information within the charts. Notably, InternVL-Chat-V1.5 experiences a decrease in its overall score and execution rate. This decline may arise from the reason that user-customized data increases the burden on the models to process different modalities of information. Overall, except for GPT-4V, the performance of the remaining LMMs is still below 60.0 in the Customized Mimic task. This indicates that, in addition to the capability to recognize data in the chart, understanding the layout of the chart and code generation are also key factors for LMMs to better accomplish the task.
Table 4: The ChartMimic leaderboard with Customized Mimic task. The best scores are in bold. We also include the code execution success rate (Exec. Rate) and model size (Params).
Model Params Exec. Rate Low-Level High-Level Overall
Text Layout Type Color Avg. GPT-4V
Proprietary
GeminiProVision - 74.0 47.2 64.9 56.0 46.4 53.6 46.8 50.2
Claude-3-opus - 88.2 65.2 82.8 48.2 39.6 58.9 54.4 56.7
GPT-4V - 93.2 77.4 89.2 78.3 65.6 77.6 72.1 74.9
Open-Weight
Qwen-VL-Chat 9.6B 6.8 2.6 3.8 3.5 2.6 3.2 2.2 2.7
LLaVA-Next-Vicuna-7B 7.1B 8.4 2.2 4.8 3.6 3.0 3.4 2.7 3.1
LLaVA-Next-Vicuna-13B 13.4B 32.2 9.3 21.5 17.4 14.2 15.6 10.8 13.2
Cogvlm2-llama3-chat-19B 19.2B 37.8 19.1 25.6 17.6 17.1 19.9 15.0 17.4
LLaVA-Next-Mistral-7B 7.6B 48.0 21.8 30.4 22.0 21.7 24.0 17.4 20.7
IDEFICS2-8B 7.6B 52.4 23.6 33.5 24.1 16.2 24.4 19.8 22.1
LLaVA-Next-Yi-34B 34.8B 59.4 33.4 45.1 26.2 24.9 32.4 17.7 25.0
InternVL-Chat-V1.5 26.0B 48.6 30.1 38.2 27.4 22.5 29.5 23.3 26.4
DeepSeek-VL-7B 7.3B 55.2 28.2 45.8 40.4 33.5 37.0 28.4 32.7
MiniCPM-Llama3-V2.5 8.4B 69.2 40.0 53.9 43.1 35.8 43.2 32.1 37.7
Phi-3-Vision 4.2B 67.2 30.9 52.5 46.6 42.0 43.0 35.7 39.4
4 Discussion
4.1 Different Complexity Levels
Figure 3: Overall scores of top-$6$ performed models at different complexity levels.
As shown in Sec. 2.3, the performance of the top-$6$ models across different complexity levels in ChartMimic is reported. With increasing difficulty, there is a consistent decline in performance across all tasks, resembling a stepped pattern. For example, in Customized Mimic task, the performance of GPT-4V at easy, medium, and hard levels are $82.8$, $69.8$, and $61.6$, respectively. Further observations across all tasks indicate that the performance decline of open-source models from easy to medium difficulty is more pronounced than from medium to hard. In contrast, proprietary models exhibit a consistent rate of decline. This suggests that the foundational capabilities of the open-source models are comparable, and the observed performance variations may largely be influenced by the quality of data. These results demonstrate that the benchmark tasks are inherently challenging and confirms the efficacy of the established difficulty levels.
4.2 Different Prompting Methods
Table 5: Results for different prompting methods on GPT-4V on Direct Mimic task.
Model Low-Level High-Level Overall
Text Layout Type Color Avg. GPT-4V
Direct 74.3 73.8 87.1 62.1 74.3 68.4 71.4
HintEnhanced 71.0 71.6 85.5 59.1 71.8 67.3 69.6
SelfRevision 71.0 70.8 84.3 59.9 71.5 65.1 68.3
Scaffold 67.4 67.5 81.7 56.3 68.2 64.4 66.3
We further explore the impact of different prompting methods. Specifically, we choose the best-performing model, GPT-4V [2], to study its performance on the Direct Mimic task. We select three representative methods for prompting: HintEnhanced, SelfRevison, and Scaffold Prompting [37]. HintEnhanced uses prompt with chain-of-thought [38], explicitly prompting the model to pay attention to important details (e.g., layout, type, text, etc). SelfRevision involves inputting the model’s own output and the rendered chart as additional information, hoping that the model can self-reflect its results and then produce better. Scaffold Prompting [37] which overlays a dot matrix within the image as visual information anchors and leverages multi-dimensional coordinates as textual positional references. Detailed introductions of these prompting methods can be found in Appendix D.
As demonstrated in Table 5, the Direct method proves to be the most straightforward and effective for GPT-4V. Other methods, to varying degrees, tend to reduce performance. Upon examining the cases, we find that the HintEnhanced method, which involves generating captions first, can introduce hallucinations that lead to errors in the subsequently generated code. Case studies suggest that although SelfRevision can identify discrepancies between output results, it often provides incorrect suggestions, resulting in subsequent errors. As for the Scaffold, its implementation can interfere with existing coordinate axis information in charts with high information density, thereby negatively impacting performance. We provide error cases of three prompting methods in Appendix D.
4.3 Correlation with Human Evaluation
Table 6: Pearson correlation coefficient between multi-level evaluation metric and human evaluation.
Metric Coefficient p-value
High-Level 0.6942 $<0.0001$
Low-Level 0.7538 $<0.0001$
To evaluate the reliability of the proposed automatic metrics, we calculate their correlation with human evaluations. Specifically, we collect 1,200 charts collected from GPT-4V using four different prompting methods (Sec. 4.2) on Direct Mimic. Each charts is evaluated by three individual evaluators, who assign scores ranging from 00 to $100$ based on the similarity with the ground-truth chart. Details about the human evaluation process are provided in Appendix E. As shown in Table 6, the Pearson correlation coefficient ($r$) indicates that both the high-level ($r=0.6942$) and low-level metric ($r=0.7538$) have a high correlation with human judgment, demonstrating the reliability of our multi-level metrics. Moreover, the low-level metric demonstrates a higher correlation with human judgments compared to the high-level metric. This is attributed to its dependence on code execution logic, enabling it to capture detailed elements.
4.4 Error Analysis
Figure 4: Error analysis of GPT-4V across four error types on the Direct Mimic task. The number in brackets indicates the count of error case. Error examples can be found in Appendix F.
We conduct a detailed error analysis of the state-of-the-art GPT-4V model to elucidate the current limitations of LMMs on ChartMimic, thereby identifying potential areas for further improvement. Our analysis identifies and category four main types of errors, detailed below.
Code-related Errors. (1) Dimension: Errors associated with data dimension, e.g., data dimension does not satisfy required conditions for operations; (2) Access: Errors occurring when accessing iterable elements out of bounds or accessing undeclared variables; (3) Parameter: Errors related to passing incorrect parameters when invoking functions. As observed in Fig. 4 (a), the majority of errors stem from Dimension, suggesting analysis and operation on data poses a challenge for GPT-4V.
Text-related Errors. (1) Detail: The text in the generated chart is largely consistent with the ground-truth chart text but contains minor discrepancies, such as mixing up “-” and “_”; (2) Missing: The generated chart omits text present in the ground-truth chart; (3) Extraneous: The generated chart contains text not found in the ground-truth chart, e.g., adding text for titles, reflecting the model’s own interpretation of the chart; As depicted in Fig. 4 (b), the majority of errors are Missing, which indicates that for GPT-4V, due to intensive information in charts, it is challenging for it to comprehend the whole scope of the charts, even the basic text recognition task.
Type-related Errors. (1) Confusion: GPT-4V misinterprets the chart type as one that appears or functions similarly, such as mistaking violin plots for box plots, and so on; (2) Missing: The generated chart omits chart types present in the ground-truth chart; (3) Extraneous: The generated chart includes chart types not present in the ground-truth chart. As demonstrated in Fig. 4 (c), although GPT-4V exhibits remarkable capability in object recognition for natural images, it still struggles with scientific charts, which contain more nuanced semantic meanings through visual logic.
Color-related Errors. (1) Similar: The colors are not the same as the ground-truth colors but appear analogous. (2) Different: The colors are entirely dissimilar to the ground-truth colors. As shown in Fig. 4 (d), though GPT-4V can not exactly recognize the accurate colors, it can identify similar ones.
To sum up, GPT-4V still faces challenges in code generation and exhibits notable visual understanding deficiencies on ChartMimic. It has difficulty accurately recognizing visual elements in figures or may hallucinate incorrect elements and struggle with complex data analysis. Combining these insights, there is a need for further improvement in both visual understanding and code generation.
5 Related Work
Large Multimodal Models. The proprietary LMMs such as GPT-4V [2], Gemini [3], and Claude-3 [28] have enabled complex multimodal interactions. Similarly, emerging open-weight LMMs such as LLaVA [33, 30], Qwen-VL [34], DeepSeek-VL [31] have contributed to the community. Despite these advancements, the effective evaluation of LMMs remains a major challenge. However, effectively evaluating LMMs remains a challenge, with open-source models performing well in benchmarks [12, 39, 40, 21, 41, 42, 43] yet falling short in practical applications [44, 45, 25]. This gap emphasizes the need for real-world-based evaluations. ChartMimic addresses this by requiring LMMs to translate complex visual information into code, testing their visual and coding capabilities.
Code Generation. Tasks such as HumanEval [6], MBPP [7], APPS [8], and DS-1000 [9] are important benchmarks in natural language processing, but the inputs for these tasks are single-modal, consisting only of text. With the emergence of LMMs, the evaluation of multimodal code generation has become more important. On top of this, MMCode [24] attempts to address this issue, but their visual inputs may be simplistic and face the problem of a single evaluation method. Design2Code [25] is relevant to our work, and they evaluate LMMs’ code generation abilities through HTML web page generation. However, their test data comes from the C4 [27] dataset, which may pose a risk of data leakage. Meanwhile, they just imitate the HTML and do not take the customized instructions; we take this into consideration. Our ChartMimic provides a new set of manually curated $1000$ data pairs, ensuring chart diversity and offering more fine-grained evaluation methods.
Chart Understanding. ChartMimic evaluates the capabilities of LMMs in grounding chart understanding into code generation. Previous works focus on chart question answering [21, 46, 39] and chart captioning [23, 22]. They assess the LMMs’ ability to understand specific data characteristics or summarize key information into text. ChartMimic advances the field by introducing a chart-to-code task, transforming the LMMs’ understanding of charts into code, which is neglected before but a realistic scenario for usage. This approach enables a comprehensive evaluation of the LMMs’ overall comprehension of charts and their ability to express this understanding in code form. Leveraging the linguistic properties of code, our benchmark introduces fine-grained metrics to assess LMMs’ chart understanding capabilities across multiple dimensions, including text, chart type, layout and color.
6 Conclusions
In this study, we develop the ChartMimic benchmark to evaluate LMMs’ proficiency capability via chart-to-code generation. ChartMimic focuses on real-world applications for data visualization, aiming to assess LMMs’ ability to harmonize a blend of cognitive capabilities, including visual understanding, code generation, and cross-modal reasoning. We propose two distinct levels of evaluation metrics (low and high level) to provide a comprehensive assessment. ChartMimic directly contributes to the understanding of progress towards artificial general intelligence, reflecting the expertise and reasoning abilities expected of skilled adults in various professional fields. Despite its comprehensive nature, ChartMimic, like any benchmark, has limitations. The manual curation process, although thorough, may introduce biases. Additionally, using scientific charts as information-intensive visual inputs to measure LMMs’ multimodal code generation capabilities, while effective, still encounters domain-specific challenges. Our evaluation metric, despite considering most elements’ similarity, does not uniformly score details of sub-icons, such as markers. We anticipate that ChartMimic will inspire the development of LMMs, advancing the pursuit of artificial general intelligence. Future research could explore various aspects, such as multimodal reasoning prompt strategies, to further reduce the gap between open-weight LMMs and proprietary ones.
Appendix
[…]
Appendix A Data Annotation
A.1 Data Annotation Principles
Diversity of Chart Types. Data visualization has become an essential tool for conveying information in various fields, and the design practices and requirements for different types of charts vary significantly[47]. Most of previous work [48, 39, 21, 22, 23] focus only on line, bar, pie charts, etc, which are commonly used within the field of computer science. However, with the increasing integration of LMMs into everyday tasks, people from all fields are starting to use generative techniques as daily assistants to enhance their design and creative processes when creating visualizations. In this light, enriching the spectrum of chart diversity is crucial for evaluating LMMs’ proficiency in multimodal chart-to-code generation.
Balance of Charts Complexity. Charts serve as visual aids for data presentation, enabling users to promptly grasp the underlying patterns and significance within the data. It is essential to adopt the appropriate chart type and complexity level to effectively convey the information [49]. Previous works have focused mainly on charts with only a single data format and low information density, which are rarely encountered in practical settings such as academic writing. Noting that LMMs like GPT-4V have already demonstrated outstanding data visualization capabilities, our focus is on using charts that are actually employed in practice, such as those from research papers, and on selecting charts with varying levels of complexity when constructing our benchmark.
Reduction of Data Leakage. Recognizing that augmenting training data is a primary method for enhancing the performance of LMMs has led researchers to more comprehensively exploit all accessible data during pre-training. However, this approach introduces the possibility of data leakage, especially when pre-training data might already include resources such as the matplotlib gallery [^2] https://matplotlib.org/ or other pre-existing datasets, potentially resulting in inaccurate evaluations. To mitigate this issue, we deliberately avoid the use of code that can be readily found online or code that could be auto-generated by large language models for chart creation in constructing our dataset, thereby reducing the probability of data leakage. To further illustrate the potential data leakage in matplotlib gallery, we select the code for $20$ charts from the matplotlib gallery. Specifically, we provide the first half of the code as a prefix for Llama3-8B [15] and let the model complete the remaining code. Then we calculate the edit distance between the generated complete code block and the ground truth. Similarly, we apply the same process to the code in Direct Mimic, calculating the edit distance between the generated results and the ground truth. We find that the edit distance for the code in the matplotlib gallery is $22.1$, while the edit distance for Direct Mimic’s code is $39.8$. This indicates that the code in Direct Mimic have a larger edit distance, which further reduces the risk of data leakage compared to the matplotlib gallery ones.
Integration of Authentic User Requirements. Users from diverse domains such as finance, healthcare, education, and engineering demonstrate unique needs and preferences for data visualization. These requirements extend beyond mere diversity; they demand the incorporation of charts capable of articulating complex and multi-dimensional data, and conforming to domain-specific aesthetic preferences[50, 49]. By aligning our dataset construction process with these real-world demands, we enable a more relevant and precise evaluation of LMMs. Adopting this approach not only reflects actual user patterns but also steers the research community in the iterative improvement of LMMs. This focused development will also lead to an enhanced user experience and increased user satisfaction, as the models become more proficient in meeting the sophisticated and varied needs of users.
A.2 Data Annotation Pipeline
General Filtering. To obtain a high-quality dataset aligned with real-world use cases and to avoid data leakage, we initially scrape figures from source files of publications on arXiv [^3] https://arxiv.org/ that hold a CC BY 4.0 license and have a publication date after February 2024. We then extract figures in PDF format that contain at least one chart, yielding approximately $174,100$ figures across various domains (such as Physics, Mathematics, Computer Science, Quantitative Biology, Quantitative Finance, Statistics, Electrical Engineering and Systems Science, Economics). We further filter the figures based on criteria of how designers select inspiring visualization examples, including clarity and visual appeals[13, 14], color schemes[13], and uniqueness of the chart within its category, resulting in a refined collection of about $15,800$ charts. It is noteworthy that during this process, some charts, which do not entirely meet the clarity and aesthetic criteria, are retained in the corpus for their unique and rare qualities, pending further modifications.
Diversity and Information Density Filtering. In this phase, we adopt a two-stage filtering process focused on the diversity of the charts and the effectiveness of information communication, primarily assessed based on the complexity of visual elements and the information density of each chart[13]. Specifically, during the first round, we maintain a data pool, categorizing chart types and subcategories. Each chart is compared against the most similar one within the corresponding category in the data pool. If significant differences in visual complexity or information density are observed, the chart is added to the pool. Otherwise, we retain the one that better fulfills the criteria for diversity and effective information communication.
After the first round of filtration, we observe a decreasing frequency of charts with increasing levels of difficulty among the remaining $15,800$ entries. Charts with simpler designs and lower information density often share greater similarity in their arrangement of visual elements, leading to a higher rate of duplicates being identified and removed. In contrast, more complex charts with higher information density are less likely to have comparable counterparts, thereby reducing duplication and enhancing the retention rate of these complex charts. This process effectively refines the complexity spectrum of our dataset, ensuring a balanced distribution.
Subsequently, we conduct the second round of filtration. We engage five experts from different fields—computer science, visual communication design, and data visualization—to independently assess the data, creating five separate data pools. We then preserve the intersection of these pools and conduct a final, collective filtration from the combined set. The selection of which charts are to be included in the final pool is made through a voting process involving all five experts. Ultimately, this approach results in a curated collection of 279 charts.
Manual Selection and Filling. In addition sourcing from arXiv, we curate chart figures from diverse platforms such as the matplotlib gallery, Stack Overflow, and plotting-related forums on Twitter and Reddit. These charts are deliberately chosen for their distinctive styles, which are not present in our arXiv dataset. To prevent data leakage, we meticulously process the data and styling of these charts, substituting existing style schemes and data with those not present in our dataset, while preserving their visual appeal.
Code and Instruction Writing. We propose to manually write codes and instructions for ChartMimic based on the collected $500$ prototype charts. To ensure the quality of the annotation results, we assemble a team of skilled Python users. These individuals are master students from computer science-related disciplines, with at least 4 years of experience using Python and extensive familiarity with matplotlib. Initially, the annotators are tasked with meticulously reproducing the $500$ prototype charts using Python and the matplotlib package, resulting in a set of $500$ (figure, code, instruction) triplets for the Direct Mimic task. Subsequently, to simulate the real-world scenario of Customized Mimic, annotators are instructed to modify the chart data in the Direct Mimic task by integrating new data from various domains. They are then required to modify the corresponding codes and instructions, leading to the $500$ (figure, code, instruction) triplets for the Customized Mimic task. Consequently, we obtain the ChartMimic benchmark, comprising $1,000$ high quality test examples.
Appendix B Chart Taxonomy
This section presents the chart taxonomy in ChartMimic. It encompasses a structure of 22 categories according to chart type characteristics and data composition. The categories comprise of:
  • 18 regular types, ordered as follows: Bar, Heatmap, Scatter, Box, Errorbar, Errorpoint, Line, Violin, Radar, Pie, Density, Graph, Quiver, Contour, Histogram, Tree, Area, and 3D charts.
  • 4 advanced types: PIP (Plot-in-Plot), Multidiff (Multiple Differences), Combination, and HR (Hard-to-Recognize).
The regular types are further divided into subcategories according to chart feature or data characteristics, whereas each chart of the advanced types represents a unique subcategory. The taxonomy showcases examples to showcase diversity of each category.
Bar: Bar chart uses rectangular bars to represent data and can be distinguished by its orientation, horizontal or vertical, with its nuanced data attributes. There are 8 subcategories of the data attributes for each orientation, consisting of 16 subcategories in total, as shown in Figure 5:
  1. 1.
    Base (single positive data set, unordered)
  2. 2.
    Sorted (data in ascending or descending sequence)
  3. 3.
    Grouped (multiple positive data sets, adjacent)
  4. 4.
    Stacked (multiple positive data sets laying atop one another)
  5. 5.
    Normalized (proportioned stacks of positive data summing to one, a Stacked variant)
  6. 6.
    Diverging (multiple stacked data sets expanding from a central axis)
  7. 7.
    With-Negative (data sets including negative values)
  8. 8.
    Reverse (exclusively negative data sets)
Figure 5: Examples of Bar chart subcategories.
Heatmap: Based on the layout and visual representations, Heatmap is categorized into 4 subcategories, each reflecting unique aspects of data presentation, as shown in Figure 6.
  1. 1.
    Base (the general layout of a typical heatmap)
  2. 2.
    Missing-Data (visualization indicating the absence of data points)
  3. 3.
    Triangle-Layout (heatmap configured in a triangular layout)
  4. 4.
    Other-Shaped (heatmap comprising elements in non-rectangular forms, such as circles)
Figure 6: Examples of Heatmap subcategories.
Scatter: Scatter plot is classified based on the dot characteristics and data distribution. The taxonomy is divided into 4 subcategories as shown in Figure 7.
  1. 1.
    Base (basic scatter plot, uniform dot size, color may vary)
  2. 2.
    Diff-Shape (different dot shapes)
  3. 3.
    Diff-Size (different dot sizes, such as a bubble chart)
  4. 4.
    Clustered (scatter plot with clear clustering)
Figure 7: Examples of Scatter plot subcategories.
Box: Box plot is characterized by the orientation of boxes and data characteristics. The orientation can be horizontal or vertical, while data characteristics include the grouping of data and the presence of missing lines. The taxonomy is divided into 3 subcategories for each orientation, consisting of 6 subcategories in total, as shown in Figure 8.
  1. 1.
    Base (single group of data, complete box shape)
  2. 2.
    Grouped (multiple groups of data, complete box shape)
  3. 3.
    Missed-Line (missing parts below the first quartile line and above the third quartile line)
Figure 8: Examples of Box chart subcategories.
Error Bar: An Error bar chart is an enhanced variant of the basic bar chart, augmented with error margins to represent the variability or uncertainty within the data. Unlike Bar chart, Error Bar chart typically does not include the categories "Normalized" and "Sorted" in the dimension of data attributes, as these are less common. Therefore, Error Bar chart is classified into 6 data attributes subcategories for each orientation, resulting in 12 subcategories in total, as shown in Figure 9.
  1. 1.
    Base (single positive data set, unordered)
  2. 2.
    Grouped (multiple positive data sets, adjacent)
  3. 3.
    Stacked (multiple positive data sets laying atop one another)
  4. 4.
    Diverging (multiple stacked data sets expanding from a central axis)
  5. 5.
    With-Negative (data sets including negative values)
  6. 6.
    Reverse (exclusively negative data sets)
Figure 9: Examples of Error Bar chart subcategories.
Error Point: Error Point chart enhances the classic Scatter chart by introducing error bars to each data point, conveying the inherent variability or uncertainty of the data. When classifying Error Point chart, a pivotal consideration lies in the orientation and symmetry of the error bars, leading us to define 5 key characteristics that govern their taxonomy. These characteristics—symmetry and asymmetry in both horizontal and vertical orientations, coupled with a composite category encompassing both directions—culminate into 5 comprehensive subcategories, as illustrated in Figure 10. Here we contend the following delineations:
  1. 1.
    Vertical-Horizontal Orientation: Distinguishing the direction of error bars, which can profoundly affect the interpretation of the data.
  2. 2.
    Symmetry-Asymmetry: Acknowledging whether error bars exhibit a mirrored consistency or an uneven distribution across data points.
Figure 10: Examples of Error Point chart subcategories.
Line: Line chart is classified based on three primary attributes: data grouping, error visualization methods, and dot-line characteristics for grouped data, as shown in Figure 11.
  1. 1.
    Data grouping: Single (individual dataset) or Grouped (multiple datasets).
  2. 2.
    Error visualization methods: Base (no error), Striped-Error (striped fill patterns), and Marker-Error (markers above and below data points).
  3. 3.
    Dot-Line Characteristics: For single dataset, No-Marker (data points without markers), Marker (data points with markers). For grouped datasets, Diff-Color (different colors for each group), Diff-Marker (different marker shapes for each group), and Diff-Line (different line styles for each group).
Figure 11: Examples of Line chart subcategories.
Violin: Violin chart is a combination of the Box chart and kernel Density chart. It provides a deep insight into the distribution of the data, indicating where individual data points fall within the overall data range. Based on the number of data groups and the shape and distribution of the violin form, the taxonomy is divided into 3 subcategories, as shown in Figure 12.
  1. 1.
    Base (standard shape, single data group)
  2. 2.
    Grouped-Symmetrical (standard shape, multiple data groups)
  3. 3.
    Grouped-Departed (half shapes joined, multiple data groups)
Figure 12: Examples of Violin chart subcategories.
Radar: Radar chart, also known as spider chart or star plot, is a graphical method of displaying multivariate data on a two-dimensional plane. It is particularly useful for showing performance metrics or skill assessments across multiple areas. Radar chart is classified based on three primary attributes: the number of data grouping, area fill, and dot-line characteristics for grouped data, as shown in Figure 13.
  1. 1.
    Data Grouping: Base (a single dataset) or Grouped (multiple datasets).
  2. 2.
    Area Fill: FillArea (areas within the radar chart filled) or NoFillArea (areas without fill to emphasize the outline).
  3. 3.
    Dot-Line Characteristics: For single dataset, NoMarker (dots without markers), Marker (dots with markers). For grouped datasets, Diff-Color (different colors for each group), Diff-Line (different line styles for each group), and Diff-Marker (different marker shapes for each group).
Figure 13: Examples of Radar chart subcategories.
Pie: Pie chart is a circular statistical graphic that divides a circle into slices to illustrate numerical proportion, whereas ring charts, also known as donut charts, utilize a hollow circle to serve a similar purpose. The presence of a ring, the number of layers, and the highlighted segment are the primary attributes in the classification of pie and ring charts, as shown in Figure 14.
  1. 1.
    Hollowness: Pie (no ring) or Ring (with a ring).
  2. 2.
    Layering: SingleLayer (single data series) or MultiLayer (multiple series or categories).
  3. 3.
    Highlighting: Base (without highlighted segments) or Explode (with one or more segments emphasized to capture viewer attention).
Figure 14: Examples of Pie and Ring chart subcategories.
Density: Density chart conveys the concentration and distribution of data within a space, often used to depict the magnitude or frequency across different areas or intervals. The orientation and data grouping are the primary attributes of the taxonomy of density plots, as shown in Figure 15.
  1. 1.
    Orientation: Vertical (y-axis as density) or Horizontal (x-axis as density).
  2. 2.
    Data Grouping: Base (single dataset) or Grouped (multiple datasets).
Figure 15: Examples of Density chart subcategories.
Graph: Graph chart commonly represents relationships and interconnected data through nodes (also known as vertices) and edges. It is widely used to depict networks, pathways, and complex inter-dependencies. The taxonomy of graph charts is based on the directionality and weight of the edges, resulting in 4 subcategories, as shown in Figure 16.
  1. 1.
    Directionality: Directed (edges have direction) or Undirected (edges have no direction).
  2. 2.
    Weight: Weighted (edges have weight) or Unweighted (edges have no weight).
Figure 16: Examples of Graph chart subcategories.
Quiver: Quiver chart, also known as vector field plot, is used to display the magnitude and direction of vectors across a two-dimensional plane. This type of chart is particularly useful in physics and engineering to represent velocity fields, gradients, or other vector-based data spatially. The taxonomy of quiver charts is based on the vector quantity and data grouping, resulting in 4 subcategories, as shown in Figure 17.
  1. 1.
    Vector Quantity: Simple (limited vectors) or Field (vector field).
  2. 2.
    Data Groups: Single (single group) or Grouped (multiple groups).
Figure 17: Examples of Quiver chart subcategories.
Contour: Contour chart, also called contour map or isoline graph, is used to represent three-dimensional data on a two-dimensional plane by plotting contour lines that connect points of equal value. This method is especially useful in fields like meteorology and geography, where it visually communicates variations in terrain elevation or changes in meteorological elements like temperature and pressure. The taxonomy of Contour chart is based on the representation of the contour lines, resulting in 3 subcategories, as shown in Figure 18.
  1. 1.
    Line (line representation)
  2. 2.
    Fill-Area (color-filled representation)
  3. 3.
    Combination (both line and color-filled)
Figure 18: Examples of Contour chart subcategories.
Histogram: Histogram chart, often referred as hist chart, is a representation of data distribution where the data is grouped into ranges or “bins” and illustrated as bars to show the frequency of data points within each bin. It is particularly useful for identifying patterns or anomalies in the data, such as skewness, peaks, or gaps in the distribution. The taxonomy of histogram charts is based on the data grouping and positioning, resulting in 3 subcategories, as shown in Figure 19.
  1. 1.
    Base (single dataset)
  2. 2.
    Overlaid (overlapping multiple datasets)
  3. 3.
    Stacked (stacked multiple datasets)
Figure 19: Examples of Histogram chart subcategories.
Treemap: Treemap displays hierarchical data through nested rectangles, where each branch of the tree is represented by a rectangle that contains smaller rectangles corresponding to sub-branches. This method allows for efficient use of space, enabling the viewer to quickly compare sizes and proportions within the hierarchy, and is especially useful for analyzing large datasets to reveal relationships and patterns. The taxonomy of treemap charts is based on the compactness and edge presence, resulting in 4 subcategories, as shown in Figure 20.
  1. 1.
    Tight-Edge (compact with border)
  2. 2.
    Tight-NoEdge (compact without border)
  3. 3.
    Loose-Edge (loose with border)
  4. 4.
    Loose-NoEdge (loose without border)
Figure 20: Examples of Treemap chart subcategories.
Area: Area chart is a graphical representation where data points are connected by line segments and the area between the line and the axis is filled with color or patterns, providing a sense of volume. It is particularly useful for visualizing the cumulative magnitude of values over time, allowing for a clear perception of trends and changes in the data series. The taxonomy of area charts is based on the presence of markers, resulting in 2 subcategories as shown in Figure 21.
  1. 1.
    Base (without markers)
  2. 2.
    Marker (with markers)
Figure 21: Examples of Area chart subcategories.
3D charts: 3D chart extends two-dimensional charting into three dimensions with spatial representations, offering an added layer of depth to represent additional data variables or to enhance visual appeal. Based on the above chart types and the additional surface representation in 3D space, we classify 3D chart into 5 subcategories, as shown in Figure 22.
  1. 1.
    Scatter (3D scatter chart)
  2. 2.
    Surface (3D surface chart)
  3. 3.
    Line (3D line chart)
  4. 4.
    Bar (3D bar chart)
  5. 5.
    Density (3D density plot)
Figure 22: Examples of 3D chart subcategories.
PIP: PIP chart insets a magnified or reduced portion of the main plot to highlight key data features, as in Figs. 23 (1) and (2). They enhance readability for complex datasets and facilitate comparative analyses, as demonstrated in Figs. 23 (3) and (4).
Figure 23: Examples of PIP charts.
Multidiff: A Multidiff chart combines at least two different chart types across multiple subplots, with each subplot presenting one type. Derived from the categories above and using diverse layouts, Multidiff charts offer numerous configurations, as shown in Fig. 24.
Figure 24: Examples of Multidiff charts.
Combination: Combination chart merges features from different chart types into one plot, offering multilayered presentation. Unlike Multidiff chart with multiple subplots for different categories, Combination chart displays multiple categories in a single plot. As in Fig. 25 (2), a scatter plot illustrates the data distribution while adjacent density plots detail the axis-specific spread. Additional examples are shown in Fig. 25.
Figure 25: Examples of Combination charts.
HR: An HR chart is one that defies the above 21 categories or is challenging to identify. HR chart is typically modified from common charts with distinctive features like custom visual arrangement or atypical markers, as shown in Fig.26.
Figure 26: Examples of HR charts.
Appendix C Details of Evaluation Metrics
In this section, we present the details of evaluation metrics, including GPT-4V score for high-level metrics and Text Score, Layout Score, Type Score and Color Score for low-level metrics. For more information, please refer to our source code.
C.1 GPT-4V Score
We employ GPT-4V [2] to assess the extend to which the generated figure corresponds to the ground-truth figure. The specific content of the prompt is presented in Fig. 27. Specifically, we input both the generated and the ground-truth figures into the GPT-4V simultaneously. Then, GPT-4V is instructed to evaluate the similarity between the two figures, taking into account six dimensions: text, layout, type, data, style, and clarity. Subsequently, GPT-4V outputs a score ranging from 0 to 100 to represent the degree of similarity between the figures.
Prompt for GPT-4V Score You are an excellent judge at evaluating visualization chart plots. The first image (reference image) is created using ground truth matplotlib code, and the second image (AI-generated image) is created using matplotlib code generated by an AI assistant. Your task is to score how well the AI-generated plot matches the ground truth plot. ### Scoring Methodology: The AI-generated image’s score is based on the following criteria, totaling a score out of 100 points: 1. Chart Types (20 points): Does the AI-generated image include all chart types present in the reference image (e.g., line charts, bar charts, etc.)? 2. Layout (10 points): Does the arrangement of subplots in the AI-generated image match the reference image (e.g., number of rows and columns)? 3. Text Content (20 points): Does the AI-generated image include all text from the reference image (e.g., titles, annotations, axis labels), excluding axis tick labels? 4. Data (20 points): How accurately do the data trends in the AI-generated image resemble those in the original image and is the number of data groups the same as in the reference image? 5. Style (20 points): Does the AI-generated image match the original in terms of colors (line colors, fill colors, etc.), marker types (point shapes, line styles, etc.), legends, grids, and other stylistic details? 6. Clarity (10 points): Is the AI-generated image clear and free of overlapping elements? ### Evaluation: Compare the two images head to head and provide a detailed assessment. Use the following format for your response: Comments: - Chart Types: \${your comment and subscore} - Layout: \${your comment and subscore} - Text Content: \${your comment and subscore} - Data: \${your comment and subscore} - Style: \${your comment and subscore} - Clarity: \${your comment and subscore} Score: \${your final score out of 100} Please use the above format to ensure the evaluation is clear and comprehensive.
Figure 27: Prompt for GPT-4V Score.
C.2 Text Score
Listing 1: An exemplary Python code for logging text information.
from matplotlib.backends.backend_pdf import RendererPdf
drawed_texts = []
def log_function(func):
def wrapper(*args, **kwargs):
global drawed_texts
text_string = args[4]
drawed_texts.append( text_string )
return func(*args, **kwargs)
return wrapper
RendererPdf.draw_text = log_function(RendererPdf.draw_text)
In order to accurately capture the textual content presented in the rendered figures, our code tracer monitors the function used to add text to the rendered PDF, logging each textual element. An exemplary Python code is provided in Listing LABEL:lst:text, where we record the text elements by adding a log wrapper to the “draw_text()” function in the matplotlib package. Subsequently, we employ this approach to extract the text information from both the generated code and the ground-truth code. The F1 score, computed between these two groups of texts, is denoted as Text Score.
C.3 Layout Score
Listing 2: An exemplary Python code for logging layout information.
def get_gridspec_layout_info(fig):
Get the layout information of a given figure.
Args:
fig (matplotlib.figure.Figure): The figure to extract layout information from.
Returns:
layout_info (list): A list of dictionaries, each containing layout information for each subplot in the figure.
layout_info = {}
for ax in fig.axes:
spec = ax.get_subplotspec()
if spec is None:
continue
gs = spec.get_gridspec()
nrows, ncols = gs.get_geometry()
row_start, row_end = spec.rowspan.start, spec.rowspan.stop - 1
col_start, col_end = spec.colspan.start, spec.colspan.stop - 1
layout_info[ax] = dict(nrows=nrows, ncols=ncols, row_start=row_start, row_end=row_end, col_start=col_start, col_end=col_end)
layout_info = list(layout_info.values())
return layout_info
layout_info = get_gridspec_layout_info(fig=plt.gcf())
The layout refers to the organization of subplots within a chart figure. In each figure implemented using matplotlib, multiple axes objects are present, each containing layout information that specifies its position within the figure. To analyze the layout, we iterate through each axis object in the figure and obtain their respective layout information. An exemplary Python code to accomplish this process is provided in Listing LABEL:lst:layout, where the position information of each axis is recorded. Subsequently, we gather the layout information from both the ground-truth code and the generated code, and compute the F1 score between the them. This metric is referred as the Layout Score.
C.4 Type Score
Listing 3: An exemplary Python code for logging type information.
from matplotlib.axes import Axes
import inspect
called_functions = {}
def log_function(func):
def wrapper(*args, **kwargs):
file_name = inspect.getfile(func)
name = file_name + "/" + func.__name__
called_functions[name] = called_functions.get(name, 0) + 1
result = func(*args, **kwargs)
return func(*args, **kwargs)
return wrapper
Axes.bar = log_function(Axes.bar)
The matplotlib package provides a variety of functions for easily generating diverse chart types, such as using “axes.bar()” to create bar charts. By monitoring the invocations of these functions, we can identify the types of charts being utilized. A successful invocation of a plot functions indicates the incorporation of a specific chart type in the final visualization. Listing LABEL:lst:charttype presents an exemplary Python code that demonstrates this approach, where we introduce a logger to the “bar()” function. Subsequently, we gather the chart types from both the generated and ground-truth code, and compute the F1 score between them. We refer to this metric as Type Score.
C.5 Color Score
Listing 4: An exemplary Python code for logging color information.
from matplotlib.axes import Axes
import inspect
drawed_colors = []
def log_function(func):
def wrapper(*args, **kwargs):
func_name = inspect.getfile(func) + "/" + func.__name__
result = func(*args, **kwargs)
for item in result:
color = item.get_facecolor()
drawed_colors.append( func_name + "--" + color )
Axes.bar = log_function(Axes.bar)
In the matplotlib package, each plot function returns a chart type instance at the end of the function invocation. These instances contain various attributes, including those related to color properties, such as facecolor, edgecolor and colormap. To assess the color attributes, we employ the code tracer that captures the color information of each chart type instance. An example Python code demonstrating this approach is provided in Listing LABEL:lst:color. Once we obtain the color attributes from both the ground-truth code and the generated code, we compare them using the F1 score metric.
It is noteworthy that through preliminary experiments, we find that using exact color matching resulted in very low color similarity, as even slight variations would result in a similarity score of zero. To address this issue, we employ the CIEDE2000 color difference formula[51], which converts the matching value between two colors from a discrete [0,1] scale to a continuous range between 0 and 1. Finally, we calculate the maximum color similarity between the sets of colors in the ground-truth code and the generated code to compute F1 score. We refer to the F1 score as Color Score.
Appendix D Generation Configurations and Prompting Methods
D.1 Generation Configurations
Following previous setups [52, 53], for open-weight models, we set the temperature $\tau=0.1$ to achieve optimal results, while for proprietary models, we set the temperature $\tau=0$ for greedy decoding. For all models, we set the maximum generation length to 4096. Additionally, we use BF16 for model inference for open-weight models. All models are inferred on a single A100 80G GPU.
D.2 Prompts
We provide prompts for Direct, HintEnhanced, SelfRevision and Scaffold Prompting in Figs. 28, 29, 30, 31 and 32. for Direct Prompting, we meticulously design a separate prompt for open-source models to achieve optimal results, as shown inFig. 28. while the prompt for proprietary Models is shown in Fig. 29.
Prompt for Direct Prompting (Open-Weight Models) You are an expert Python developer who specializes in writing matplotlib code based on a given picture. I found a very nice picture in a STEM paper, but there is no corresponding source code available. I need your help to generate the Python code that can reproduce the picture based on the picture I provide. Please note that it is necessary to use figsize=({width} (width), {height} (height)) to set the image size to match the original size. Additionally, I will not provide you with the actual data in the image, so you have to extract the actual data by yourself and based on the extracted data to reproduce the image. Ensure that the code you provide can be executed directly without requiring me to add additional variables. Now, please give me the matplotlib code that reproduces the picture below.
Figure 28: Prompt for Direct Prompting (Open-Weight Models). {text} in blue font represents placeholders, which varies according to different test examples.
Prompt for Direct Prompting (Proprietary Models) You are an expert Python developer who specializes in writing matplotlib code based on a given picture. I found a very nice picture in a STEM paper, but there is no corresponding source code available. I need your help to generate the Python code that can reproduce the picture based on the picture I provide. Note that it is necessary to use figsize=({width} (width), {height} (height)) to set the image size to match the original size. Now, please give me the matplotlib code that reproduces the picture below.
Figure 29: Prompt for Direct Prompting (Proprietary Models). {text} in blue font represents placeholders, which varies according to different test examples.
Prompt for HintEnhanced Prompting You are an expert Python developer who specializes in writing matplotlib code based on a given picture. I found a very nice picture in a STEM paper, but there is no corresponding source code available. I need your help to generate the Python code that can reproduce the picture based on the picture I provide. To ensure accuracy and detail in your recreation, begin with a comprehensive analysis of the figure to develop an elaborate caption. This caption should cover, but not be limited to, the following aspects: 1. Layout Analysis: e.g., identify the picture’s composition, noting the presence and arrangement of any subplots. 2. Chart Type Identification: e.g., determine how many charts within a subplot. Are they independent, or do they share a common axis? 3. Data Analysis: e.g., summarize the data trend or pattern. 4. Additional Features: e.g.,identify any supplementary elements such as legends, colormaps, tick labels, or text annotations that contribute to the figure’s clarity or aesthetic appeal. Now, given the picture below, please first output your comprehensive caption and then use the caption to assist yourself to generate matplotlib code that reproduces the picture. Note that it is necessary to use figsize=({width} (width), {height} (height)) to set the image size to match the original size.
Figure 30: Prompt for HintEnhanced Prompting. {text} in blue font represents placeholders, which varies according to different test examples.
Prompt for SelfRevision Prompting You are an expert Python developer who specializes in writing matplotlib code based on a given picture. I have a code for implementing the reference picture as follows: Now, I have the Python matplotlib code for implementing the reference picture as follows: ```python {python_function} ``` The rendered picture of the code is: Now, please compare whether the renderer picture is the same as the reference picture. The difference may cover, but not be limited to, the following aspects: 1. Chart Types: Does the AI-generated image include all chart types present in the reference image (e.g., line charts, bar charts, etc.)? 2. Layout: Does the arrangement of subplots in the AI-generated image match the reference image (e.g., number of rows and columns)? 3. Text Content: Does the AI-generated image include all text from the reference image (e.g., titles, annotations, axis labels), excluding axis tick labels? 4. Data: How accurately do the data trends in the AI-generated image resemble those in the original image and is the number of data groups the same as in the reference image? 5. Style: Does the AI-generated image match the original in terms of colors (line colors, fill colors, etc.), marker types (point shapes, line styles, etc.), legends, grids, and other stylistic details? 6. Clarity: Is the AI-generated image clear and free of overlapping elements? - If the generated picture matches the reference, please output the original implementation code. - If there are discrepancies, first list the specific differences between the two pictures. Then, modify the existing code to address these differences, ensuring the revised code is capable of reproducing the reference picture. Finally, output the revised code. Note that it is necessary to use figsize=({width} (width), {height} (height)) to set the image size to match the original size.
Figure 31: Prompt for SelfRevision Prompting. {text} in blue font represents placeholders, which varies according to different test examples.
Prompt for Scaffold Prompting You are an expert Python developer who specializes in writing matplotlib code based on a given picture. I found a very nice picture in a STEM paper, but there is no corresponding source code available. I need your help to generate the Python code that can reproduce the picture based on the picture I provide. I will provide you with two images. The first image is the original picture. The second image is the picture overlaid with a dot matrix of a shape of {dot_matrix_height} * {dot_matrix_width} to help you with your task, and each dot is labeled with two-dimensional coordinates (x,y). Within each column, the x-coordinate increases from top to bottom, and within each row, the y-coordinate increases from left to right. Please first use this dot matrix as reference anchors to generate the description of the picture (e.g., between dot A and dot B is something) and generate matplotlib code that reproduces the picture. Note that it is necessary to use figsize=({width} (width), {height} (height)) to set the image size to match the original size.
Figure 32: Prompt for Scaffold Prompting. {text} in blue font represents placeholders, which varies according to different test examples.
D.3 Error Cases of Different Prompting Methods
We provide error cases of HintEnhanced, SelfRevision and Scaffold Prompting in Figs. 33, 34 and 35. For an analysis of error cases related to Direct Prompting, please refer to Appendix F.
Figure 33: An error case of HintEnhanced Prompting. The text highlighted in blue is GPT-4V’s incorrect understanding regarding provided figures. The remaining part of response is omitted with “……”.
Figure 34: An error case of SelfRevision Prompting. The text highlighted in blue is GPT-4V’s incorrect understanding regarding provided figures. The remaining part of response is omitted with “……”.
Figure 35: An error case of Scaffold Prompting. The text highlighted in blue is GPT-4V’s incorrect understanding regarding provided figures. The remaining part of response is omitted with “……”.
Appendix E Correlation with Human Evaluation
Figure 36: A screenshot of the human evaluation questionnaire.
In order to evaluate the reliability of the proposed automatic metrics, we calculate their correlation with human evaluations. Specifically, a selection of $300$ test examples from the Direct Mimic task is utilized. Subsequently, we gather the results generated by GPT-4V using four different prompting methods (Sec. 4.2) on these test examples. This process yields $1,200$ figures to be assessed.
Each figure is independently assessed by three evaluators through a questionnaire, who assign scores on a scale of 00 to $100$ based on the similarity between the generated figures and the ground-truth figures. The final score for each figure is calculated as the mean of scores given by the three evaluators. Our evaluators comprise volunteer graduate students holding bachelor’s degrees in computer science. To facilitate their assessments, we provide a scoring rubric in the questionnaire, which closely aligns with the low-level and high-level metric evaluations. The scoring criteria encompass the following dimensions: chart type, layout, textual content, data, style, and clarity. A screenshot of the final questionnaire is available in Fig. 36.
Upon collecting the human evaluation results, we calculate the Pearson correlation coefficient to ascertain the correlation between our automatic metrics and human evaluation.
Appendix F Cases of Error Analysis
We provide cases of text-related, type-related and color-related errors in LABEL:fig:error_1, LABEL:fig:error_2, LABEL:fig:error_3, LABEL:fig:error_4, LABEL:fig:error_5 and LABEL:fig:error_6. These cases encompass various error types mentioned in Sec. 4.4.
Appendix G Ethics and Societal Impact
We envision ChartMimic as a comprehensive suite of benchmarks designed to guide researchers in assessing the capabilities of their LMMs. By providing an extensive evaluation framework, ChartMimic aims to foster the development of foundation models for the community, offering insights into various aspects of LMMs, such as visual understanding, code generation, and cross-modal reasoning.
When constructing prototype charts, we use figures from papers on arXiv, which are licensed under CC BY 4.0. During the manual annotation of the code, we implement strict ethical controls to ensure that the information conveyed in the code does not contain any biased content and does not infringe on any private data.

### v2

@@
 Abstract
-We introduce a new benchmark, ChartMimic, aimed at assessing the visually-grounded code generation capabilities of large multimodal models (LMMs).
+We introduce a new benchmark, ChartMimic, aimed at assessing the visually grounded code generation capabilities of large multimodal models (LMMs).
 ChartMimic utilizes information-intensive visual charts and textual instructions as inputs, requiring LMMs to generate the corresponding code for chart rendering.
-ChartMimic includes $1,000$ human-curated (figure, instruction, code) triplets, which represent the authentic chart use cases found in scientific papers across various domains (e.g., Physics, Computer Science, Economics, etc).
-These charts span $18$ regular types and $4$ advanced types, diversifying into $191$ subcategories.
+ChartMimic includes $4,800$ human-curated (figure, instruction, code) triplets, which represent the authentic chart use cases found in scientific papers across various domains (e.g., Physics, Computer Science, Economics, etc).
+These charts span $18$ regular types and $4$ advanced types, diversifying into $201$ subcategories.
 Furthermore, we propose multi-level evaluation metrics to provide an automatic and thorough assessment of the output code and the rendered charts.
 Unlike existing code generation benchmarks, ChartMimic places emphasis on evaluating LMMs’ capacity to harmonize a blend of cognitive capabilities, encompassing visual understanding, code generation, and cross-modal reasoning.
-The evaluation of $3$ proprietary models and $11$ open-weight models highlights the substantial challenges posed by ChartMimic.
-Even the advanced GPT-4V, Claude-3-opus only achieve an average score of $73.2$ and $53.7$, respectively, indicating significant room for improvement.
+The evaluation of $3$ proprietary models and $14$ open-weight models highlights the substantial challenges posed by ChartMimic.
+Even the advanced GPT-4o, InternVL2-Llama3-76B only achieved an average score across Direct Mimic and Customized Mimic tasks of $82.2$ and $61.6$, respectively, indicating significant room for improvement.
 We anticipate that ChartMimic will inspire the development of LMMs, advancing the pursuit of artificial general intelligence.
 1 Introduction
-Code generation [1] is a rather demanding task that requires advanced abstract thinking and logical reasoning, reflecting the unique intelligence of human beings.
-Recently, progresses in artificial general intelligence (AGI) have demonstrated the potential of large foundation models [2, 3] in solving these tasks that were once the exclusive domain of human abilities [4, 5].
-However, existing code generation benchmarks [6, 7, 8, 9] solely use text as input, while humans receive information from multiple modalities when coding [10, 11], such real-life scenario has not been fully explored.
+Code generation (Sun et al., 2024) is a rather demanding task that requires advanced abstract thinking and logical reasoning, reflecting the unique intelligence of human beings.
+Recently, advances in artificial general intelligence (AGI) have demonstrated the potential of large foundation models (Google, 2023; OpenAI, 2024; Anthropic, 2024; AI@Meta, 2024) to solve the tasks that are once the exclusive domain of human abilities (Achiam et al., 2023; Zhu et al., 2024).
+However, existing code generation benchmarks (Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021; Lai et al., 2023) solely use text as input, while humans receive information from multiple modalities when coding (Liang et al., 2023; Fan et al., 2024).
+Such real-life scenarios have yet to be fully explored.
 Taking a common scene in Fig.
@@
 2), a multimodal code generation benchmark.
-ChartMimic is characterized by its (1) information-intensive visual inputs, (2) diverse chart types and (3) multi-level evaluation metrics.
+ChartMimic is characterized by its (1) information-intensive visual inputs, (2) diverse chart types, and (3) multi-level evaluation metrics.
 Specifically, compared to natural images, scientific charts convey nuanced semantic meanings through intricate visual logic, thereby exhibiting higher information density.
 Based on this, we define two tasks, Direct Mimic and Customized Mimic (Sec.
-2.1), utilizing charts and textual instructions as inputs.
+2.1), which utilize charts and textual instructions as inputs.
 These tasks challenge LMMs to generate the corresponding code for a given chart or to incorporate new data specified in the instructions, respectively.
-Through the collection of academic documents and scientific papers, we identify $22$ commonly used chart types and $191$ subcategories.
-Subsequently, we manually annotate a total of $1,000$ (figure, instruction, code) triplets for these types (Sec.
+Through the collection of academic documents and scientific papers, we identify $22$ commonly used chart types and $201$ subcategories.
+Subsequently, we manually annotate a total of $4,800$ (figure, instruction, code) triplets for these types (Sec.
 2.2).
@@
 2.4).
-We conduct examination of $14$ LMMs on ChartMimic (Sec.
-3.2), including $3$ proprietary models and $11$ open-weight models.
-We observe that while several open-weight models can match the performance of proprietary models such as GPT-4V on public leaderboards [12], a significant performance gap still persists on ChartMimic.
-Specifically, the best open-weight model, Phi-3-Vision, significantly lags behind GPT-4V, with only half the performance on two tasks, indicating substantial room for improvement.
-Additionally, correlation analysis (Sec.
-4.3) demonstrates the high correlation between our automatic metrics and human judgment.
+We conduct the examination of $17$ LMMs on ChartMimic (Sec.
+3.2), including $3$ proprietary models and $14$ open-weight models across parameter sizes from 2.2B to 76.0B.
+We observe that while several open-weight models can match the performance of proprietary models such as GPT-4o on public leaderboards (OpenCompass, 2023), a significant performance gap still persists on ChartMimic.
+Specifically, the best open-weight model, InternVL2-Llama3-76B, lags behind GPT-4o, with an average score gap of $20.6$ on two tasks, indicating substantial room for improvement in the open-source community.
+Our analysis of prompting methods (Sec.
+4.2) reveals that GPT-4o can improve itself through self-reflection, which is a key manifestation of System 2 reasoning (Sloman, 1996; Kumar et al., 2024).
+This discovery highlights the vital role that the System 2 reasoning plays in LMMs when tackling the complex challenges presented by ChartMimic.
+Meanwhile, Correlation analysis (Sec.
+4.3) demonstrates a high correlation between our automatic metrics and human evaluation, validating the effectiveness of these metrics.
 Further error analysis (Sec.
-4.4) reveals that hallucination notably hinders LMMs’ performance on ChartMimic, involving the insertion of non-existent text in the ground-truth figures and confusion of similar chart types.
+4.4) reveals that hallucinations notably hinder the performance of LLMs on ChartMimic, as they lead to the insertion of non-existent text into ground-truth figures and confusion between similar types of charts.
 Figure 1: The real-world example.
@@
 These models serve as assistants that enhance the comprehension and presentation of data in scholarly communications.
-We envision ChartMimic as a suite of benchmarks designed to guide researchers in understanding the capabilities of their LMMs.
-By providing a comprehensive evaluation framework, ChartMimic aims to facilitate the growth of foundation models for the community, offering insights into LMM’s various aspects, such as visual understanding, code generation, and cross-modal reasoning.
+We envision ChartMimic as a comprehensive suite of benchmarks designed to guide researchers in understanding their LMMs’ capabilities.
+By providing a comprehensive evaluation framework, ChartMimic aims to facilitate the growth of foundation models for the community, offering insights into various aspects such as visual understanding, code generation, and cross-modal reasoning.
 2 The ChartMimic Benchmark
-In this section, we first introduce the definition of two tasks involved in ChartMimic, and then delineate the data curation process.
-Subsequently, we conduct a quantitative analysis to demonstrate the quality and diversity of ChartMimic and compare it with existing related benchmarks.
+In this section, we first introduce the definition of two tasks involved in ChartMimic (Sec.
+2.1), and then delineate the data curation process (Sec.
+2.2).
+Subsequently, we conduct a quantitative analysis to assess the quality and diversity of ChartMimic (Sec.
+2.3), establish evaluation metrics (Sec.
+2.4), and compare it with existing related benchmarks (Sec.
+2.5).
 2.1 Task Definition
-The capability of LMMs to generate code that can accurately render a specific chart demonstrates their proficiency in visual understanding, code generation and cross-modal reasoning.
+LMMs’ ability to generate chart-rendering code demonstrates their visual understanding, coding, and cross-modal reasoning skills.
 Specifically, given the chart $X$ and the instruction $I$, the LMM $f$ is expected to output the code $C$ that satisfies the requirements outlined in the instruction:
@@
 Figure 2: The pipeline of ChartMimic.
-We provide $1,000$ human-curated (figure, instruction, code) triplets.
+We provide $4,800$ human-curated (figure, instruction, code) triplets.
 We use ChartMimic to evaluate LMMs’ proficiency in the multimodal chart-to-code generation, resulting in both high-level and low-level evaluation results.
-Direct Mimic.
-The LMM is requested to directly generate code that can reproduce the provided chart.
-This task evaluates the LMM’s ability to comprehensively understand and reason about the chart to generate executable code.
+Direct Mimic. The LMMs are tasked to directly generate code that can reproduce the provided chart, thereby assessing their visual comprehension and reasoning capabilities.
 Customized Mimic.
-With additional customized data provided in the instruction, this task requires the LMM to generate code that can render a new chart encoding the customized data while maintaining the aesthetic and design of the provided chart.
-This task requires the LMM to understand both the chart and customized data to integrate them effectively.
+The LMMs are requested to generate code for a new chart that incorporates customized data provided in the instruction while preserving the original chart’s aesthetic and design, assessing their ability to integrate visual and textual information.
 After obtaining the code generated by LMMs, we execute it to render the corresponding chart and subsequently compare its similarity with the ground-truth chart.
@@
 2.2 Data Curation Process
-ChartMimic distinguishes itself through $4$ fundamental considerations: (1) diversity of chart types, (2) balance of charts complexity, (3) reduction of data leakage, (4) integration of authentic user requirements.
-We keep on these principles to complete the data curation for ChartMimic through a four-step pipeline.
-We provide an overview here and more details in Appendix A.
+ChartMimic distinguishes itself through $4$ fundamental considerations: (1) diversity of chart types, (2) balance of chart complexity, (3) reduction of data leakage, (4) integration of authentic user requirements.
+We keep on these four principles to complete the data curation for ChartMimic through a five-step pipeline.
+Here, we provide an overview here and more details in Appendix B.
 General Filtering.
-We scrape figures from source files of publications on arXiv [^1] https://arxiv.org/ that hold a CC BY 4.0 license and have a publication date after February 2024.
-We extract figures in PDF format, yielding approximately $174,100$ figures across various domains (e.g., Physics, Computer Science, Economics, etc).
-We then filter the figures based on the criteria of how designers select inspiring visualization examples, including clarity and visual appeals [13, 14], color schemes [13], and uniqueness of the chart within its category, resulting in a refined collection of about $15,800$ figures.
+We scrape figures from source files of publications on arXiv [^1] https://arxiv.org/ that hold a CC BY 4.0 license with a publication date after February 2024 with PDF format.
+This yields approximately $174,100$ figures across various domains (e.g., Physics, Computer Science, Economics, etc).
+We then filter the figures based on the criteria of how designers select inspiring visualization examples, including clarity and visual appeal, color schemes (Bako et al., 2022; Quispel et al., 2018), and uniqueness of the chart within its category, resulting in a refined collection of about $15,800$ figures.
 Diversity and Information Density Filtering.
-We first initialize a data pool, creating a taxonomy of chart types and subcategories and comparing each chart against the most similar one within the corresponding category in the data pool.
-Should significant differences in visual complexity or information density [13] be found, the chart is added to the pool.
-Otherwise, we retain the one that better fulfills the criteria for diversity and effective information communication.
-Ultimately, this approach yields a curated collection of $279$ charts from the generally filtered $15,800$ ones.
+The filtering process involves two stages.
+In the first stage, we establish a data pool and categorize chart types.
+Charts with significant differences in complexity or information density are added to ensure diversity and effective communication (Bako et al., 2022).
+In the second stage, five experts from various fields independently evaluate the data, creating separate selection pools.
+We preserve the intersection of their selections and finalize the set through a voting process.
+This meticulous approach refines our collection to $279$ charts.
 Manual Selection and Filling.
@@
 To mitigate the risk of data leakage, we rigorously process the data and color styling of these charts, replacing existing color schemes with those unseen in our data pool while maintaining their aesthetic appeal.
-Consequently, we obtain $500$ prototype charts for ChartMimic.
+Consequently, we obtain $600$ prototype charts for ChartMimic.
 Code and Instruction Writing.
-We propose to manually write codes and instructions for ChartMimic based on the collected 500 prototype charts.
-Initially, the annotators are tasked with meticulously reproducing the 500 prototype charts using Python code, resulting in a set of 500 (figure, code, instruction) triplets for the Direct Mimic task.
+We propose to manually write codes and instructions for ChartMimic based on the collected $600$ prototype charts.
+Initially, annotators are tasked with meticulously reproducing the $600$ prototype charts using Python code, resulting in a set of $600$ (figure, code, instruction) triplets for the Direct Mimic task.
+Although other coding languages such as JavaScript and R can be used to create charts, current LMMs perform poorly when doing chart-to-code using these other languages (Sun et al., 2024).
+Therefore, we focus on Python code generation in the current version of ChartMimic.
 Subsequently, to simulate the scenario of Customized Mimic, annotators are instructed to modify the chart data in the Direct Mimic task by integrating new data from various domains.
-They are then required to modify the corresponding codes and instructions, leading to the 500 (figure, code, instruction) triplets for the Customized Mimic task.
-Consequently, we obtain the ChartMimic benchmark, comprising $1,000$ high quality test examples.
-Table 1: Statistics of ChartMimic.
-Each one in Combination, HR, Muiltidiff, and PIP is an independent subcategory.
-We measure code length in terms of tokens, utilizing the Llama3 tokenizer [15].
-In the level count, “A/B/C” denotes the number of chart for “easy/medium/hard” level, respectively.
+They are then required to modify the corresponding code and instructions, leading to the $600$ (figure, code, instruction) triplets for the Customized Mimic task.
+Consequently, we establish the ChartMimic benchmark, comprising $1,200$ high-quality seed data.
+Table 1: Statistics of ChartMimic. We measure code length in terms of tokens, utilizing the Llama3 tokenizer. In the level count, “A/B/C” denotes the number of “easy/medium/hard” level, respectively.
 Type Bar Line ErrorBar Heatmap Box Scatters Hist Radar 3D Pie ErrorPoint Violin
-Count 100 80 30 30 25 25 20 20 15 15 10 10
+Count 320 280 120 120 100 100 80 80 80 80 80 80
 Subcategories 16 8 12 4 6 4 3 6 5 8 5 3
-Code Length (AVG.) 689.3 805.7 651.9 664.6 659.9 617.2 507.3 772.2 668.7 385.1 523.7 912.0
-Code Length (STD.) 289.0 272.2 146.1 303.0 248.9 287.6 161.2 136.4 168.9 123.1 228.9 298.7
-Level Count 53 / 40 / 7 74 / 6 / 0 17 / 13 / 0 0 / 19 / 11 15 / 10 / 0 20 / 5 / 0 13 / 7 / 0 13 / 7 / 0 2 / 8 / 5 10 / 5 / 0 7 / 3 / 0 4 / 6 / 0
+Code Length (AVG.) 689.6 794.0 681.2 685.8 689.0 655.0 529.6 779.8 655.4 418.4 624.3 975.6
+Code Length (STD.) 237.8 244.4 144.7 258.7 228.2 253.0 147.1 144.3 241.4 99.5 197.7 252.3
+Level Count 176 / 120 / 24 256 / 24 / 0 68 / 52 / 0 0 / 76 / 44 60 / 40 / 0 80 / 20 / 0 52 / 28 / 0 52 / 28 / 0 8 / 48 / 24 52 / 28 / 0 44 / 28 / 8 32 / 44 / 4
 Type Area Contour Density Graph Quiver Treemap Combination HR Muiltidiff PIP Total
-Count 5 5 5 5 5 5 30 25 25 10 500
-Subcategories 2 3 4 4 4 4 30 25 25 10 101+(90)
-Code Length (AVG.) 832.8 352.8 431.8 310.8 845.0 302.2 641.9 689.9 788.1 958.9 682.9
-Code Length (STD.) 177.5 136.7 71.9 61.7 607.0 34.7 159.0 301.5 331.0 285.9 285.1
-Level Count 4 / 1 / 0 0 / 0 / 5 3 / 2 / 0 5 / 0 / 0 0 / 2 / 3 5 / 0 / 0 3 / 19 / 8 1 / 4 / 20 0 / 12 / 13 0 / 0 / 10 249 / 169 / 82
-2.3 Data Statistics
-Notably, the (figure,code,instruction) triplets for both Direct and Customized Mimic tasks share the same figure.
-Therefore, we detail the data statistics for the $500$ triplets in the Direct Mimic task.
+Count 80 80 80 80 80 80 120 100 100 80 2400
+Subcategories 2 3 4 4 4 4 30 25 25 20 201(101 + 100)
+Code Length (AVG.) 774.4 489.4 540.0 564.5 893.4 342.2 697.4 718.9 798.2 1083.9 696.0
+Code Length (STD.) 161.8 87.8 104.7 117.5 631.0 36.3 163.6 265.5 271.2 290.1 278.4
+Level Count 52 / 28 / 0 0 / 28 / 52 44 / 32 / 4 56 / 24 / 0 0 / 52 / 28 52 / 28 / 0 12 / 76 / 32 4 / 16 / 80 0 / 48 / 52 0 / 0 / 80 1100 / 868 / 432
+Data Augmentation.
+Following the development of seed triplets, we initiate a process of manual data augmentation.
+Annotators are tasked with altering various elements of each seed triplet, including data, color schemes, mark styles, etc., to produce augmented triplets.
+For each seed triplet, we create $3$ additional augmented triplets.
+This process enhances our dataset, yielding a total of $4,800$ triplets that reflect a wide range of realistic and practical chart use cases.
+2.3 Data Statistics and Division
+The (figure, code, instruction) triplets for both Direct and Customized Mimic tasks share the same figure.
+Therefore, we detail the data statistics for the $2,400$ triplets in the Direct Mimic task.
 As depicted on the left side of Fig.
 2, ChartMimic encompasses a total of $22$ categories, with $18$ types of regular charts and $4$ types of advanced charts.
-For the $18$ regular chart types, we identify $101$ subcategories, with the definitions and examples of each subcategory provided in Appendix B.
+For the $18$ regular chart types, we identify $101$ subcategories, with the definitions and examples of each subcategory provided in Appendix D.
 The advanced chart types such as Plot-in-Plot (PIP), Multidiff, and Combination are distinct forms of amalgamating multiple chart sets into a singular chart.
 Given the variety of their internal combination elements, each of these can be treated as a unique subcategory.
-Meanwhile, the Hard-to-Recognize (HR) category encapsulates unclassifiable charts, with each chart being considered a category unto itself.
-When we factor in the additional $90$ subcategories represented by these advanced chart types, ChartMimic encompasses a total of $191$ subcategories.
+Meanwhile, the Hard-to-Recognize (HR) category encapsulates unclassifiable charts, with each chart in the seed data being considered a category unto itself.
+When we factor in the additional $100$ subcategories represented by these advanced chart types, ChartMimic encompasses a total of $201$ subcategories.
 This extensive diversity underscores the comprehensive nature of our benchmark.
-Additionally, we employ Llama3 [15] tokenizer to measure the code length.
-The results are shown in Table 1, which shows an average code token length of $635.98$ with a standard deviation of $253.06$.
-Moreover, we conduct a manual complexity assessment for each chart, classifying them into three levels: easy, medium, and hard.
-The distribution of charts according to complexity levels is as follows: $249$ for easy, $169$ for medium, and $82$ for hard.
+We employ Llama3 (AI@Meta, 2024) tokenizer to measure the code length.
+As shown in Tab.
+1, ChartMimic has an average code token length of $696.0$ with a standard deviation of $278.4$.
+Additionally, we manually categorize charts into three complexity levels: easy (1,100), medium (868), and hard (432).
+The detailed categorization criteria and assessment methodology are thoroughly documented in Appendix B.
+We further divide the $4,800$ examples of ChartMimic into two subsets: test and testmini set.
+The test set comprises $3,600$ examples, while the testmini set is composed of $1,200$ examples.
+The testmini set is designed for rapid model development validation.
+Our partitioning strategy ensures each chart type is proportionally represented, preserving a distribution in the testmini set that closely aligns with the test set.
+Detailed comparative experimental results, discussed in Appendix C, demonstrate the consistency across two subsets.
+Unless otherwise stated, we report results on the testmini set.
 2.4 Evaluation Metrics
 For tasks within ChartMimic, an appropriate evaluation necessitates comparing the visual similarity between the generated and ground-truth figures.
-To achieve automatic evaluation, we propose multi-level metrics (i.e., low-level and high-level), which assess the similarity at different granularities.
-We compute the average score between the low-level and high-level scores, which range from 0 to 100, as the overall score.
+To achieve this, we propose multi-level metrics (i.e., high-level and low-level) to assess the similarity at different granularities.
+Specifically, the high-level metric encompasses the GPT-4o Score, and the low-level metric encompasses Text, Layout, Type and Color Score.
+We compute the average score between the high-level and low-level scores, ranging from 0 to 100, as the overall score.
 The illustration of the multi-level metrics is depicted in Fig.
 2.
-GPT-4V Score.
-Following the successful utilization of large foundation models like GPT-4 for automated evaluation in both natural language processing [16, 17, 18] and computer vision [19, 20], we adopt the GPT-4V score as our high-level metric.
-Specifically, we input both the ground-truth chart and the generated chart to GPT-4V, and instruct GPT-4V to output a high-level similarity score between 00 and $100$.
-The detailed prompt template can be found in Appendix C.
+GPT-4o Score.
+Following the successful use of large foundation models for evaluation in both natural language processing (Zheng et al., 2024; Li et al., 2023; Dubois et al., 2024) and computer vision (Zhang et al., 2023; Yang et al., 2024; Wu et al., 2024b), we adopt GPT-4o score as our high-level metric.
+Specifically, we input both the ground-truth figure and the generated figure into GPT-4o, and instruct it to output a high-level similarity score ranging 00 to $100$.
+Although CLIP Score (Radford et al., 2021) is widely used for assessing image similarity, in our preliminary experiments, it has struggled to distinguish variations in types and other critical elements in charts, resulting in a low correlation with human evaluation results.
+Therefore, we use only GPT-4o Score as our high-level evaluation metric.
+The detailed description of GPT-4o Score can be found in Appendix E.
 Table 2: A comparison of our proposed ChartMimic to other benchmarks. “I” and “NL” indicate “Image” and “Natural Language” respectively.
@@
 Chart Understanding Benchmarks
-ChartQA [21] Human Curated 3 10K I+NL NL Accuracy
-Chart-to-Text [22] Crawl 6 44K I+NL NL Match-based
-ChartSumm [23] Human Curated 3 84K I+NL NL Match-based
+ChartQA (Masry et al., 2022) Human Curated 3 10K I+NL NL Accuracy
+Chart-to-Text (Kantharaj et al., 2022) Crawl 6 44K I+NL NL Match-based
+ChartSumm (Rahman et al., 2023) Human Curated 3 84K I+NL NL Match-based
+CharArXiv (Wang et al., 2024b) Human Curated 18 93K I+NL NL GPT-4 Score
 Code Generation Benchmarks
-HumanEval [6] Human Curated - 164 Code Code Pass Rate
-MBPP [7] Human Curated - 500 NL+Code Code Pass Rate
-MMCode [24] Crawl - 263 I+NL Code Pass Rate
-MatPlotBench [20] Human Curated 13 100 NL Code GPT-4V Score
-Design2Code [25] Crawl HTML 484 I+NL Code Multi-Level
-ChartMimic (Ours) Human Curated 22 1K I+NL Code Multi-Level
-In addition to high-level similarity, evaluating low-level elements between generated and ground-truth figures can help better analyze the behaviors of LMMs.
-Therefore, we propose evaluating four key low-level elements: text, layout, type and color.
-Extracting them from figures is a challenging task, as existing extraction models often fall short in terms of accuracy [26].
-Considering that figures are rendered based on the code, we design a code tracer to monitor the execution process of the ground-truth code and generated code.
+HumanEval (Chen et al., 2021) Human Curated - 164 Code Code Pass Rate
+MBPP (Austin et al., 2021) Human Curated - 500 NL+Code Code Pass Rate
+MMCode (Li et al., 2024b) Crawl - 263 I+NL Code Pass Rate
+MatPlotBench (Yang et al., 2024) Human Curated 13 100 NL Code GPT-4 Score
+Plot2Code (Wu et al., 2024a) Crawl 15 132 I+NL Code Multi-Level
+Design2Code (Si et al., 2024) Crawl HTML 484 I+NL Code Multi-Level
+ChartMimic (Ours) Human Curated 22 4,800 I+NL Code Multi-Level
+In addition to high-level similarity, evaluating the similarity among low-level elements between generated and ground-truth figures can provide a more fine-grained analysis of LMMs.
+Therefore, we propose to evaluate four key low-level elements in charts (Savva et al., 2011; Poco & Heer, 2017): text, layout, type and color.
+Extracting them from figures is a challenging task, as existing extraction models often fall short in terms of accuracy (Meng et al., 2024).
+Considering that figures are rendered based on the code, we design a code tracer to monitor the execution processes of the ground-truth code and generated code.
 The code tracer records the text, layout, type and color information.
@@
 And we average text, layout, type and color scores to get the low-level score.
-The methodology for obtaining these elements is briefly introduced as follows, with detailed descriptions provided in Appendix C.
+The methodology for obtaining these elements is briefly introduced as follows, with detailed descriptions provided in Appendix E.
 Text Score. During the code execution process, for the function responsible for adding text elements to the rendered figures, the code tracer monitors it and records each text element parameter.
@@
 Color Score. For each plot function, it will return the chart type instance at the end of the function invocation. The code tracer accesses the color attributes of these chart type instances.
+It is important to note that code execution success rate is a standard metric for code generation tasks (Sun et al., 2024).
+We have implicitly incorporated this aspect into our high-level and low-level scores.
+Specifically, if the code fails to execute successfully, both the low-level and high-level scores are assigned a value of 0.
+Therefore, we do not separately weight it into the overall score.
 2.5 Comparisons with Existing Benchmarks
-To further distinguish the difference between ChartMimic and other existing ones, we elaborate the benchmark details in Table 2.
-From the chart understanding perspective, the prior benchmarks [21, 22, 23] are heavily focused on simple questions about the data in the charts and lack of both difficulty and diversity in chart types.
-In the code generation aspect, HumanEval [6], MBPP [7] and MatPlotBench [20] only consider tasks with text inputs, which may not meet requirements in the era of LMMs.
-Recently, MMCode [24] attempted to create a benchmark for multimodal code generation, but the vision inputs for their task are still overly simple and only have a single pass rate evaluation metric.
-Design2Code [25] is the most similar one to ours, which focuses on automated webpage generation.
-Although they use multi-level evaluation metrics like ours, their test HTMLs all come from the C4 dataset [27], which may pose a risk of data leakage.
+To further distinguish the difference between ChartMimic and other existing ones, we elaborate the benchmark details in Tab.
+2.
+From the chart understanding perspective, the prior benchmarks (Masry et al., 2022; Kantharaj et al., 2022; Rahman et al., 2023; Wang et al., 2024b) are only focused on questions about the data in the charts without necessitating the advanced cognitive capabilities of LMMs, which include visual understanding, code generation, and cross-modal reasoning.
+In the code generation aspect, HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021) and MatPlotBench (Yang et al., 2024) only consider tasks with text inputs, which may not meet requirements in the era of LMMs.
+Recently, MMCode (Li et al., 2024b) attempted to create a benchmark for multimodal code generation, but the vision inputs for their task are still overly simple and only have a single pass rate evaluation metric.
+Design2Code (Si et al., 2024) and Plot2Code (Wu et al., 2024a) are the most similar ones to ours.
+Although they use multi-level evaluation metrics like ours, their test is crawled directly from the internet or existing datasets, which may pose a risk of data leakage.
 Our ChartMimic benchmark gives complex scientific charts as inputs, demanding capabilities on grounding visual understanding into code generation, and provides multi-level evaluation metrics.
@@
 3.1 Baseline Setup
-We benchmark $14$ widely utilized proprietary and open-source models currently available in the field.
-For proprietary models, we consider $3$ representative models: GPT-4V [2], Claude-3-opus [28] and GeminiProVision [3].
-For the open-weight models, we choose $11$ competitive models.
-We list them in increasing order of total parameter size from 4.2B to 34.8B, which are: Phi-3-Vision [29], LLaVA-Next-Vicuna-7B [30], DeepSeek-VL-7B [31], LLaVA-Next-Mistral-7B [30], IDEFICS2-8B [32], MiniCPM-Llama3-V 2.5 [33], Qwen-VL-Chat [34], LLaVA-Next-Vicuna-13B [30], Cogvlm2-llama3-chat-19B [35], InternVL-Chat-V1.5 [36], LLaVA-Next-Yi-34B [30].
-We start with the direct prompting method, which provides the reference chart with direct instructions.
-The specific instructions and models’ decoding setups can be found in Appendix D.
-3.2 Main Results
+We benchmark $17$ widely utilized proprietary and open-source models currently available in the field.
+For proprietary models, we consider $3$ representative models: GPT-4o (OpenAI, 2024), Claude-3-opus (Anthropic, 2024) and GeminiProVision (Google, 2023).
+For the open-weight models, we choose $14$ competitive models with total parameter size from 2.2B to 76.0B: InternVL2(2B, 4B, 8B, 26B, 76B) (Chen et al., 2023), Qwen2-VL(2B, 7B) (Wang et al., 2024a), Phi-3-Vision (phi, 2024), DeepSeek-VL-7B (Lu et al., 2024), LLaVA-Next(7B, 34B) (Li et al., 2024a), IDEFICS2-8B (Laurençon et al., 2024), MiniCPM-Llama3-V2.5 (Xu et al., 2024) and Cogvlm2-llama3-chat-19B (Wang et al., 2023a).
+We start with direct prompting, which provides the reference chart with direct instructions.
+The specific instructions and model configurations can be found in Appendix F.
 Table 3: The ChartMimic leaderboard with Direct Mimic task. The best scores are in bold. We also include the code execution success rate (Exec. Rate) and model size (Params).
 Model Params Exec. Rate Low-Level High-Level Overall
-Text Layout Type Color Avg. GPT-4V
+Text Layout Type Color Avg. GPT-4o
 Proprietary
-GeminiProVision - 64.2 36.9 58.6 45.1 39.6 45.0 38.1 41.6
-Claude-3-opus - 86.4 61.8 79.0 45.3 37.9 56.0 45.4 50.7
-GPT-4V - 91.4 74.3 87.1 73.8 62.1 74.3 68.4 71.4
+GeminiProVision - 68.2 52.6 64.2 51.3 47.1 53.8 53.3 53.6
+Claude-3-opus - 83.3 66.8 83.1 49.9 42.1 60.5 60.1 60.3
+GPT-4o - 93.2 81.5 89.8 77.3 67.2 79.0 83.5 81.2
 Open-Weight
-Qwen-VL-Chat 9.6B 4.0 0.5 2.8 1.1 0.8 1.3 0.9 1.1
-LLaVA-Next-Vicuna-7B 7.1B 6.2 0.8 2.1 1.4 1.6 1.5 1.6 1.6
-LLaVA-Next-Vicuna-13B 13.4B 11.6 1.9 5.7 4.0 4.3 4.0 2.8 3.4
-IDEFICS2-8B 7.6B 42.4 5.7 28.4 11.9 10.9 14.2 8.8 11.5
-LLaVA-Next-Yi-34B 34.8B 45.6 11.4 26.7 18.2 16.6 18.2 12.7 15.4
-LLaVA-Next-Mistral-7B 7.6B 57.6 11.3 31.5 22.3 21.9 21.8 14.8 18.3
-DeepSeek-VL-7B 7.3B 39.0 13.8 26.4 23.9 19.0 20.8 16.8 18.8
-Cogvlm2-llama3-chat-19B 19.2B 51.4 20.6 31.1 21.7 20.0 23.3 17.4 20.4
-InternVL-Chat-V1.5 26.0B 60.0 26.9 42.1 26.1 24.9 30.0 23.4 26.7
-MiniCPM-Llama3-V2.5 8.4B 58.2 22.5 40.1 31.4 26.6 30.2 24.2 27.2
-Phi-3-Vision 4.2B 66.0 36.4 45.0 39.0 34.3 38.7 25.1 31.9
-We present the main results of 14 LMMs on ChartMimic. Table 3 and Table 4 show the results on the Direct Mimic and Customized Mimic task, respectively. The key findings are as follows:
-GPT-4V performs best among proprietary models, while Phi-3-Vision excels among open-weight models.
-In proprietary models, GPT-4V achieves an overall score of $71.4$ in Direct Mimic and $72.33$ in Customized Mimic, surpassing the second-place model Claude-3-opus by $21.3$ and $21.1$, respectively.
-Within open-weight models, Phi-3-Vision, despite its 4.2B parameters, reaches the overall score of $31.9$ in Direct Mimic and $40.18$ in Customized Mimic, which outperforms LLaVA-Next-Yi-34B ($15.4$ in Direct Mimic and $20.2$ in Customized Mimic).
+IDEFICS2-8B 7.6B 49.0 6.2 33.1 9.2 9.0 14.4 17.6 16.0
+DeepSeek-VL-7B 7.3B 41.3 15.3 26.6 19.7 14.5 19.0 20.4 19.7
+LLaVA-Next-Yi-34B 34.8B 50.2 15.9 29.6 17.6 15.2 19.6 20.6 20.1
+LLaVA-Next-Mistral-7B 7.6B 59.7 14.0 31.1 19.8 17.8 20.7 21.3 21.0
+Qwen2-VL-2B 2.6B 47.0 20.1 29.5 21.3 17.9 22.2 23.4 22.8
+Cogvlm2-llama3-chat-19B 19.2B 50.5 21.3 31.8 18.4 17.0 22.1 24.5 23.3
+InternVL2-2B 2.2B 52.5 23.6 35.8 16.0 15.4 22.7 24.2 23.5
+Qwen2-VL-7B 8.2B 67.0 26.4 51.0 31.0 23.3 32.9 35.0 34.0
+InternVL2-4B 4.2B 66.2 34.7 51.7 25.2 23.6 33.8 38.4 36.1
+InternVL2-8B 8.1B 61.8 31.5 51.1 28.6 26.2 34.4 38.9 36.6
+MiniCPM-Llama3-V-2.5 8.4B 80.3 30.7 49.6 38.6 27.6 36.6 42.1 39.4
+Phi-3-Vision-128K 4.2B 66.7 37.5 49.6 37.4 29.8 38.6 41.0 39.8
+InternVL2-26B 26.0B 69.3 39.2 58.7 35.9 31.8 41.4 47.4 44.4
+InternVL2-Llama3-76B 76.0B 83.2 54.1 74.5 49.2 41.5 54.8 62.2 58.5
+3.2 Main Results
+We present the main results of $17$ LMMs on ChartMimic. Tab. 3 and Tab. 4 show the results on the Direct Mimic and Customized Mimic task, respectively. The key findings are as follows:
+GPT-4o performs best among proprietary models, while InternVL2-Llama3-76B excels among open-weight models.
+In proprietary models, GPT-4o achieves an overall score of $81.2$ in Direct Mimic and $83.2$ in Customized Mimic.
+Within open-weight models, InternVL2-Llama3-76B, reaches the overall score of $58.5$ in Direct Mimic and $64.7$ in Customized Mimic, which have compareble perfromance with proprietary model like Claude-3-opus ($60.3$ in Direct Mimic and $65.4$ in Customized Mimic).
+Meanwhile, Phi-3-Vision-128K despite its 4.2B parameters, reaches the overall score of $39.8$ in Direct Mimic and $42.1$ in Customized Mimic, which outperforms LLaVA-Next-Yi-34B ($20.1$ in Direct Mimic and $35.3$ in Customized Mimic).
 This indicates that even models with fewer parameters can achieve decent performance through a refined training process.
 A performance gap persists between open-weight LMMs and proprietary ones.
-Although several open-weight LMMs can exhibit performance comparable to GPT-4V across various benchmarks [12], even the best-performing Phi-3-Vision falls short of achieving half the performance of GPT-4V on both Direct Mimic and Customized Mimic.
+Although several open-weight LMMs can exhibit performance comparable to GPT-4o across various benchmarks (OpenCompass, 2023), even the best-performing InternVL2-Llama3-76B falls short of achieving the performance of GPT-4o both Direct Mimic and Customized Mimic.
 This apparent performance disparity demonstrates the challenging nature of our ChartMimic benchmark for current open-weight LMMs.
-Moreover, open-weight LMMs even exhibit notable deficiencies in generating executable code, with a majority of them showing an execution rate below $60\%$.
-These findings emphasize that there is still a considerable scope for the open-source community to enhance LMMs’ capabilities in terms of complex visual understanding, code generation and cross-modal reasoning.
+Moreover, open-weight LMMs even exhibit notable deficiencies in generating executable code, with a majority of them showing an execution rate below $75\%$.
+These findings highlight that there is still a considerable scope for the open-source community to enhance LMMs’ capabilities regarding complex visual understanding, code generation and cross-modal reasoning.
 When given additional user-customized data, LMMs exhibit performance improvements.
@@
 This enhancement can be attributed to the provision of user-customized data, which alleviates the burden on LMMs to recognize textual and data information within the charts.
-Notably, InternVL-Chat-V1.5 experiences a decrease in its overall score and execution rate.
+Notably, Cogvlm2-llama3-chat-19B and Qwen2-VL-2B experienced a decrease in their overall score and execution rate.
 This decline may arise from the reason that user-customized data increases the burden on the models to process different modalities of information.
-Overall, except for GPT-4V, the performance of the remaining LMMs is still below 60.0 in the Customized Mimic task.
+Overall, except for GPT-4o, the performance of the remaining LMMs is still below $65.0$ in the Customized Mimic task.
 This indicates that, in addition to the capability to recognize data in the chart, understanding the layout of the chart and code generation are also key factors for LMMs to better accomplish the task.
@@
 Model Params Exec. Rate Low-Level High-Level Overall
-Text Layout Type Color Avg. GPT-4V
+Text Layout Type Color Avg. GPT-4o
 Proprietary
-GeminiProVision - 74.0 47.2 64.9 56.0 46.4 53.6 46.8 50.2
-Claude-3-opus - 88.2 65.2 82.8 48.2 39.6 58.9 54.4 56.7
-GPT-4V - 93.2 77.4 89.2 78.3 65.6 77.6 72.1 74.9
+GeminiProVision - 76.2 52.2 70.9 56.0 49.4 57.1 59.6 58.4
+Claude-3-opus - 88.2 75.2 86.8 54.1 44.3 65.1 65.7 65.4
+GPT-4o - 96.5 88.5 92.9 79.2 67.6 82.1 84.3 83.2
 Open-Weight
-Qwen-VL-Chat 9.6B 6.8 2.6 3.8 3.5 2.6 3.2 2.2 2.7
-LLaVA-Next-Vicuna-7B 7.1B 8.4 2.2 4.8 3.6 3.0 3.4 2.7 3.1
-LLaVA-Next-Vicuna-13B 13.4B 32.2 9.3 21.5 17.4 14.2 15.6 10.8 13.2
-Cogvlm2-llama3-chat-19B 19.2B 37.8 19.1 25.6 17.6 17.1 19.9 15.0 17.4
-LLaVA-Next-Mistral-7B 7.6B 48.0 21.8 30.4 22.0 21.7 24.0 17.4 20.7
-IDEFICS2-8B 7.6B 52.4 23.6 33.5 24.1 16.2 24.4 19.8 22.1
-LLaVA-Next-Yi-34B 34.8B 59.4 33.4 45.1 26.2 24.9 32.4 17.7 25.0
-InternVL-Chat-V1.5 26.0B 48.6 30.1 38.2 27.4 22.5 29.5 23.3 26.4
-DeepSeek-VL-7B 7.3B 55.2 28.2 45.8 40.4 33.5 37.0 28.4 32.7
-MiniCPM-Llama3-V2.5 8.4B 69.2 40.0 53.9 43.1 35.8 43.2 32.1 37.7
-Phi-3-Vision 4.2B 67.2 30.9 52.5 46.6 42.0 43.0 35.7 39.4
+Qwen2-VL-2B 2.6B 35.8 17.4 23.9 19.7 16.5 19.4 21.4 20.4
+Cogvlm2-llama3-chat-19B 19.2B 38.7 19.0 27.9 16.5 15.7 19.8 21.6 20.7
+LLaVA-Next-Mistral-7B 7.6B 49.0 20.0 32.0 22.6 19.9 23.6 24.7 24.2
+IDEFICS2-8B 7.6B 49.2 21.6 32.2 18.1 12.2 21.0 27.3 24.2
+InternVL2-2B 2.2B 49.3 22.2 35.4 20.0 18.1 23.9 27.8 25.9
+LLaVA-Next-Yi-34B 34.8B 64.2 28.7 44.8 32.9 27.7 33.5 37.1 35.3
+DeepSeek-VL-7B 7.3B 59.3 27.5 47.5 36.8 31.5 35.8 39.3 37.6
+Phi-3-Vision-128K 4.2B 67.8 29.7 52.5 42.3 36.5 40.3 44.0 42.1
+InternVL2-4B 4.2B 74.0 41.3 55.6 39.6 33.1 42.4 47.8 45.1
+Qwen2-VL-7B 8.2B 73.3 41.0 56.3 43.5 34.2 43.8 47.8 45.8
+InternVL2-8B 8.1B 73.0 43.1 54.4 39.9 35.4 43.2 48.9 46.1
+MiniCPM-Llama3-V-2.5 8.4B 78.7 40.8 58.0 44.8 33.2 44.2 51.5 47.9
+InternVL2-26B 26.0B 73.7 43.9 62.3 43.5 34.3 46.0 51.1 48.6
+InternVL2-Llama3-76B 76.0B 89.8 57.8 79.0 63.5 50.5 62.7 66.7 64.7
 4 Discussion
 4.1 Different Complexity Levels
-Figure 3: Overall scores of top-$6$ performed models at different complexity levels.
-As shown in Sec.
-2.3, the performance of the top-$6$ models across different complexity levels in ChartMimic is reported.
-With increasing difficulty, there is a consistent decline in performance across all tasks, resembling a stepped pattern.
-For example, in Customized Mimic task, the performance of GPT-4V at easy, medium, and hard levels are $82.8$, $69.8$, and $61.6$, respectively.
-Further observations across all tasks indicate that the performance decline of open-source models from easy to medium difficulty is more pronounced than from medium to hard.
-In contrast, proprietary models exhibit a consistent rate of decline.
-This suggests that the foundational capabilities of the open-source models are comparable, and the observed performance variations may largely be influenced by the quality of data.
-These results demonstrate that the benchmark tasks are inherently challenging and confirms the efficacy of the established difficulty levels.
+Figure 3: Overall scores of models at different complexity levels.
+We report the performance of the top-$6$ models across different complexity levels for ChartMimic in  Fig.
+3.
+There is a consistent decline in performance across all tasks as the difficulty increases.
+For example, in Direct Mimic task, the performance of GPT-4o at easy, medium, and hard levels are $86.5$, $77.7$, and $74.8$, respectively.
+These results demonstrate that ChartMimic are inherently challenging and confirm the efficacy of the established difficulty levels.
+Additionally, for weak LMMs such as MiniCPM-Llama3-V-2.5, providing customized data at the Easy Level can help it perform tasks better (from $46.8$ to $59.7$).
+However, at the Hard Level, even when given customized data, their performance do not show improvement (from $31.9$ to $31.5$).
+This indicates that for Easy Level, complex reasoning abilities are not required, and providing data is sufficient to complete the task well.
+However, for Hard Level charts, LMMs need complex code generation capabilities grounded on visual understanding.
+In this case, even when given data, if the LMMs cannot perform the corresponding code reasoning, they still cannot complete the task.
+Moreover, when given addtional data, if the data is too long and complex, the LLM may still become confused, leading to a certain decrease in performance.
 4.2 Different Prompting Methods
-Table 5: Results for different prompting methods on GPT-4V on Direct Mimic task.
-Model Low-Level High-Level Overall
-Text Layout Type Color Avg. GPT-4V
-Direct 74.3 73.8 87.1 62.1 74.3 68.4 71.4
-HintEnhanced 71.0 71.6 85.5 59.1 71.8 67.3 69.6
-SelfRevision 71.0 70.8 84.3 59.9 71.5 65.1 68.3
-Scaffold 67.4 67.5 81.7 56.3 68.2 64.4 66.3
-We further explore the impact of different prompting methods.
-Specifically, we choose the best-performing model, GPT-4V [2], to study its performance on the Direct Mimic task.
-We select three representative methods for prompting: HintEnhanced, SelfRevison, and Scaffold Prompting [37].
-HintEnhanced uses prompt with chain-of-thought [38], explicitly prompting the model to pay attention to important details (e.g., layout, type, text, etc).
-SelfRevision involves inputting the model’s own output and the rendered chart as additional information, hoping that the model can self-reflect its results and then produce better.
-Scaffold Prompting [37] which overlays a dot matrix within the image as visual information anchors and leverages multi-dimensional coordinates as textual positional references.
-Detailed introductions of these prompting methods can be found in Appendix D.
-As demonstrated in Table 5, the Direct method proves to be the most straightforward and effective for GPT-4V.
-Other methods, to varying degrees, tend to reduce performance.
+Table 5: Results for different prompting methods for GPT-4o and different size of InternVL2 LMMs on Direct Mimic task. The best scores for each model are in bold.
+Model Params Method Exec. Rate Low-Level High-Level Overall
+Text Layout Type Color Avg. GPT-4o
+GPT-4o - Direct 93.2 81.5 89.8 77.3 67.2 79.0 83.5 81.2
+HintEnhanced 92.0 80.9 89.7 78.2 67.9 79.2 83.0 81.1
+SelfReflection 95.2 86.6 93.7 81.1 68.2 82.4 87.1 84.8
+Scaffold 91.0 81.3 90.6 75.1 63.9 77.7 80.3 79.0
+InternVL2-Llama3-76B 76.0B Direct 83.2 54.1 74.5 49.2 41.5 54.8 62.2 58.5
+HintEnhanced 78.0 54.3 71.7 50.2 41.6 54.5 60.3 57.4
+SelfReflection 82.5 57.0 76.2 56.4 45.0 58.7 63.6 61.1
+Scaffold 76.7 52.1 71.4 49.1 40.5 53.3 60.1 56.7
+InternVL2-26B 26.0B Direct 69.3 39.2 58.7 35.9 31.8 41.4 47.4 44.4
+HintEnhanced 66.8 40.4 54.8 36.6 33.0 41.2 47.0 44.1
+Self 69.0 39.5 57.5 35.6 30.5 40.8 47.1 43.9
+Scaffold 64.3 36.0 53.5 35.4 30.3 38.8 44.0 41.4
+InternVL2-8B 8.1B Direct 61.8 31.5 51.1 28.6 26.2 34.4 38.9 36.6
+HintEnhanced 58.7 30.1 38.3 30.1 27.5 31.5 35.9 33.7
+SelfReflection 56.8 27.4 39.8 25.1 22.9 28.8 33.4 31.1
+Scaffold 61.8 22.8 43.5 24.8 20.2 27.8 32.0 29.9
+InternVL2-2B 2.1B Direct 52.5 23.6 35.8 16.0 15.4 22.7 24.2 23.5
+HintEnhanced 41.5 17.3 19.4 14.3 12.9 16.0 18.5 17.2
+SelfReflection 35.5 16.3 22.2 12.8 14.8 16.5 17.3 16.9
+Scaffold 38.0 6.5 17.8 3.2 2.2 7.4 15.6 11.5
+We further examine the impact of different prompting methods on the performance of ChartMimic benchmark.
+Specifically, we choose the GPT-4o and InternVL2 series LMMs (2B, 8B, 26B, 76B) to study their performance on the Direct Mimic task.
+We select three representative prompting methods: HintEnhanced, SelfReflection, and Scaffold Prompting.
+HintEnhanced uses prompt with chain-of-thought (Wei et al., 2022), explicitly prompting the LMMs to pay attention to important details (e.g., layout, type, text, etc).
+SelfReflection (Shinn et al., 2024) involves inputting the LMMs’ own output and the corresponding rendered chart as additional information, instructing the LMM’ to self-reflect their output.
+Scaffold Prompting (Lei et al., 2024) overlays a dot matrix within the figure as visual information anchors and leverages multi-dimensional coordinates as textual positional references.
+We detail the experimental setups in Appendix F.
+As shown in Tab.
+5, SelfReflection enables GPT-4o and InternVL2-Llama3-76B to reflect on and correct their outputs, demonstrating notable improvements over Direct Prompting.
+This underscore self-reflection as a key manifestation of System 2 reasoning (Sloman, 1996; Kumar et al., 2024).
+For GPT-4o, SelfReflection enhances Text, Layout, and Type Scores, though Color Score remains unchanged due to persistent challenges in fine color discrimination.
+However, LMMs with less developed reasoning capabilities (InternVL2-26B, 8B, and 2B) show no improvement or even decline with self-reflection.
+This indicates that only LMMs with substantial reasoning capabilities can effectively engage in self-reflection and result optimization, underscoring the critical role of System 2 reasoning in addressing ChartMimic’s challenges.
+Regarding HintEnhanced and Scaffold methods, they do not enhance the performance of models and reduce performance to varying degrees.
 Upon examining the cases, we find that the HintEnhanced method, which involves generating captions first, can introduce hallucinations that lead to errors in the subsequently generated code.
-Case studies suggest that although SelfRevision can identify discrepancies between output results, it often provides incorrect suggestions, resulting in subsequent errors.
-As for the Scaffold, its implementation can interfere with existing coordinate axis information in charts with high information density, thereby negatively impacting performance.
-We provide error cases of three prompting methods in Appendix D.
+As for the Scaffold method, introduction of an additional dot matrix can interfere with existing coordinate axis information in charts with high information density, thereby negatively impacting performance.
+These negative effects intensify as LMMs’ reasoning capabilities decrease, further highlighting the importance of advanced reasoning capabilities in handling complex prompting strategies.
+We provide case studies of three prompting methods in Appendix F.
 4.3 Correlation with Human Evaluation
@@
 Metric Coefficient p-value
-High-Level 0.6942 $<0.0001$
-Low-Level 0.7538 $<0.0001$
-To evaluate the reliability of the proposed automatic metrics, we calculate their correlation with human evaluations.
-Specifically, we collect 1,200 charts collected from GPT-4V using four different prompting methods (Sec.
+High-Level 0.7041 $<0.0001$
+Low-Level 0.7681 $<0.0001$
+To evaluate the reliability of the proposed multi-level metrics, we calculate their correlation with human evaluations.
+Specifically, we collect $1,200$ charts collected from GPT-4o using four different prompting methods (Sec.
 4.2) on Direct Mimic.
 Each charts is evaluated by three individual evaluators, who assign scores ranging from 00 to $100$ based on the similarity with the ground-truth chart.
-Details about the human evaluation process are provided in Appendix E.
-As shown in Table 6, the Pearson correlation coefficient ($r$) indicates that both the high-level ($r=0.6942$) and low-level metric ($r=0.7538$) have a high correlation with human judgment, demonstrating the reliability of our multi-level metrics.
+Details about the human evaluation process are provided in Appendix G.
+As shown in Tab.
+6, the Pearson correlation coefficient ($r$) indicates that both the high-level ($r=0.7041$) and low-level metric ($r=0.7681$) have a high correlation with human judgment, demonstrating the reliability of our multi-level metrics.
 Moreover, the low-level metric demonstrates a higher correlation with human judgments compared to the high-level metric.
@@
 4.4 Error Analysis
-Figure 4: Error analysis of GPT-4V across four error types on the Direct Mimic task. The number in brackets indicates the count of error case. Error examples can be found in Appendix F.
-We conduct a detailed error analysis of the state-of-the-art GPT-4V model to elucidate the current limitations of LMMs on ChartMimic, thereby identifying potential areas for further improvement.
+Figure 4: Error analysis of GPT-4o across four error types on the Direct Mimic task. The number in brackets indicates the count of error cases. Error examples can be found in Appendix Appendix I.
+We conduct a detailed error analysis of the state-of-the-art GPT-4o model to elucidate the current limitations of LMMs on ChartMimic, thereby identifying potential areas for further improvement.
 Our analysis identifies and category four main types of errors, detailed below.
@@
 As observed in Fig.
-4 (a), the majority of errors stem from Dimension, suggesting analysis and operation on data poses a challenge for GPT-4V.
+4 (a), the majority of errors stem from Dimension, suggesting analysis and operation on data poses a challenge for GPT-4o.
 Text-related Errors.
 (1) Detail: The text in the generated chart is largely consistent with the ground-truth chart text but contains minor discrepancies, such as mixing up “-” and “_”; (2) Missing: The generated chart omits text present in the ground-truth chart; (3) Extraneous: The generated chart contains text not found in the ground-truth chart, e.g., adding text for titles, reflecting the model’s own interpretation of the chart; As depicted in Fig.
-4 (b), the majority of errors are Missing, which indicates that for GPT-4V, due to intensive information in charts, it is challenging for it to comprehend the whole scope of the charts, even the basic text recognition task.
+4 (b), the majority of errors are Missing and Extraneous, which indicates that for GPT-4o, due to intensive information in charts, it is challenging for it to comprehend the whole scope of the charts, even the basic text recognition task.
 Type-related Errors.
-(1) Confusion: GPT-4V misinterprets the chart type as one that appears or functions similarly, such as mistaking violin plots for box plots, and so on; (2) Missing: The generated chart omits chart types present in the ground-truth chart; (3) Extraneous: The generated chart includes chart types not present in the ground-truth chart.
+(1) Confusion: GPT-4o misinterprets the chart type as one that appears or functions similarly, such as mistaking violin plots for box plots, and so on; (2) Missing: The generated chart omits chart types present in the ground-truth chart; (3) Extraneous: The generated chart includes chart types not present in the ground-truth chart.
 As demonstrated in Fig.
-4 (c), although GPT-4V exhibits remarkable capability in object recognition for natural images, it still struggles with scientific charts, which contain more nuanced semantic meanings through visual logic.
+4 (c), although GPT-4o exhibits remarkable capability in object recognition for natural images, it still struggles with scientific charts, which contain more nuanced semantic meanings through visual logic.
 Color-related Errors.
@@
 As shown in Fig.
-4 (d), though GPT-4V can not exactly recognize the accurate colors, it can identify similar ones.
-To sum up, GPT-4V still faces challenges in code generation and exhibits notable visual understanding deficiencies on ChartMimic.
+4 (d), though GPT-4o can not exactly recognize the accurate colors, it can identify similar ones.
+To sum up, GPT-4o still faces challenges in code generation and exhibits notable visual understanding deficiencies on ChartMimic.
 It has difficulty accurately recognizing visual elements in figures or may hallucinate incorrect elements and struggle with complex data analysis.
@@
 Large Multimodal Models.
-The proprietary LMMs such as GPT-4V [2], Gemini [3], and Claude-3 [28] have enabled complex multimodal interactions.
-Similarly, emerging open-weight LMMs such as LLaVA [33, 30], Qwen-VL [34], DeepSeek-VL [31] have contributed to the community.
+The proprietary LMMs such as GPT-4o (OpenAI, 2024), Gemini (Google, 2023), and Claude-3 (Anthropic, 2024) have enabled complex multimodal interactions.
+Similarly, emerging open-weight LMMs such as LLaVA (Xu et al., 2024; Li et al., 2024a), InternVL (Chen et al., 2023), Qwen-VL (Bai et al., 2023), DeepSeek-VL (Lu et al., 2024) have contributed to the community.
 Despite these advancements, the effective evaluation of LMMs remains a major challenge.
-However, effectively evaluating LMMs remains a challenge, with open-source models performing well in benchmarks [12, 39, 40, 21, 41, 42, 43] yet falling short in practical applications [44, 45, 25].
-This gap emphasizes the need for real-world-based evaluations.
+However, effectively evaluating LMMs remains a challenge, with open-source models performing well in benchmarks (Liu et al., 2023; Chen et al., 2024; Yue et al., 2023; Lu et al., 2023; Luo et al., 2025) yet falling short in practical applications (Xie et al., 2024; Koh et al., 2024; Si et al., 2024; Ji et al., 2024).
+This gap emphasizes the need for real-world-based evaluations that reflect authentic use cases (Ma et al., 2024a; b).
 ChartMimic addresses this by requiring LMMs to translate complex visual information into code, testing their visual and coding capabilities.
 Code Generation.
-Tasks such as HumanEval [6], MBPP [7], APPS [8], and DS-1000 [9] are important benchmarks in natural language processing, but the inputs for these tasks are single-modal, consisting only of text.
-With the emergence of LMMs, the evaluation of multimodal code generation has become more important.
-On top of this, MMCode [24] attempts to address this issue, but their visual inputs may be simplistic and face the problem of a single evaluation method.
-Design2Code [25] is relevant to our work, and they evaluate LMMs’ code generation abilities through HTML web page generation.
-However, their test data comes from the C4 [27] dataset, which may pose a risk of data leakage.
+Tasks such as HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), APPS (Hendrycks et al., 2021), and DS-1000 (Lai et al., 2023) are important benchmarks in natural language processing, but the inputs for these tasks are single-modal, consisting only of text, limiting their scope.
+With the emergence of LMMs, the evaluation of multimodal code generation has become increasingly critical for assessing real-world capabilities (Li et al., 2024b; Si et al., 2024; Wu et al., 2024a; Ge et al., 2025).
+For example, MMCode (Li et al., 2024b) attempts to address this issue, but their visual inputs may be simplistic and face the problem of a single evaluation method.
+Design2Code (Si et al., 2024) evaluates LMMs’ code generation abilities through HTML web page generation.
+However, their test data comes from the C4 (Raffel et al., 2020) dataset, which may pose a risk of data leakage.
 Meanwhile, they just imitate the HTML and do not take the customized instructions; we take this into consideration.
-Our ChartMimic provides a new set of manually curated $1000$ data pairs, ensuring chart diversity and offering more fine-grained evaluation methods.
+Recently, Plot2Code (Wu et al., 2024a) undertakes work similar to ours, aiming to measure models’ code generation abilities through chart-to-code generation.
+Similarly, their approach of directly scraping data from the matplotlib gallery poses a risk of data leakage.
+Our ChartMimic provides a new set of manually curated $4,800$ data pairs, ensuring chart diversity and offering more fine-grained evaluation methods.
 Chart Understanding.
-ChartMimic evaluates the capabilities of LMMs in grounding chart understanding into code generation.
-Previous works focus on chart question answering [21, 46, 39] and chart captioning [23, 22].
+ChartMimic evaluates the capabilities of LMMs in grounding chart understanding into code generation, bridging visual and programmatic domains.
+Previous works focus on chart question answering (Masry et al., 2022; Methani et al., 2020; Xu et al., 2023; Wang et al., 2024b; Li et al., 2024c; Zeng et al., 2024) and chart captioning (Rahman et al., 2023; Kantharaj et al., 2022).
 They assess the LMMs’ ability to understand specific data characteristics or summarize key information into text.
-ChartMimic advances the field by introducing a chart-to-code task, transforming the LMMs’ understanding of charts into code, which is neglected before but a realistic scenario for usage.
+ChartMimic advances the field by introducing a chart-to-code task, transforming the LMMs’ understanding of charts into code, which is neglected before but a realistic scenario for practical, real-world usage.
 This approach enables a comprehensive evaluation of the LMMs’ overall comprehension of charts and their ability to express this understanding in code form.
@@
 Future research could explore various aspects, such as multimodal reasoning prompt strategies, to further reduce the gap between open-weight LMMs and proprietary ones.
+Acknowledgments
+This work was partly supported by the National Key Research and Development Program of China (No. 2024YFB2808903) and the Shenzhen Science and Technology Program JSGG20220831110203007).
 Appendix
@@
 […]
-Appendix A Data Annotation
-A.1 Data Annotation Principles
+Appendix A Author Contributions
+Yujiu Yang served as the principal advisor for this project. Cheng Yang and Chufan Shi acted as the student project leads.
+Data Annotation.
+Yaxin Liu and Bo Shui co-led the data annotation process (Appendix B).
+Chufan Shi, Cheng Yang, Mohan Jing and Linran Xu helped to contribute to various aspects of this process.
+Xiaomei Nie provided expert guidance on data visualization and conducted quality control.
+Code Implementation. Cheng Yang and Chufan Shi co-implemented the codebase for ChartMimic.
+Experiments. Cheng Yang and Chufan Shi jointly conducted the evaluation of LMMs on ChartMimic. Yaxin Liu and Bo Shui jointly led the human evaluation.
+Paper Writing.
+Cheng Yang and Chufan Shi finished the main manuscript.
+Junjie Wang and Yujiu Yang provided in-depth reviews and revisions.
+Yaxin Liu and Bo Shui created all visualizations presented in the paper and contributed to the Data Annotation (Appendix B) and Chart Taxonomy (Appendix D).
+Deng Cai, Xinyu Zhu, Siheng Li, Yuxiang Zhang and Gongye Liu contributed through proofreading, discussions and feedback.
+Appendix B Data Annotation
+B.1 Data Annotation Principles
 Diversity of Chart Types.
-Data visualization has become an essential tool for conveying information in various fields, and the design practices and requirements for different types of charts vary significantly[47].
-Most of previous work [48, 39, 21, 22, 23] focus only on line, bar, pie charts, etc, which are commonly used within the field of computer science.
+Data visualization has become an essential tool for conveying information in various fields, and the design practices and requirements for different types of charts vary significantly(Parsons, 2021).
+Most of previous work (Han et al., 2023; Xu et al., 2023; Masry et al., 2022; Kantharaj et al., 2022; Rahman et al., 2023) focus only on line, bar, pie charts, etc, which are commonly used within the field of computer science.
 However, with the increasing integration of LMMs into everyday tasks, people from all fields are starting to use generative techniques as daily assistants to enhance their design and creative processes when creating visualizations.
@@
 Charts serve as visual aids for data presentation, enabling users to promptly grasp the underlying patterns and significance within the data.
-It is essential to adopt the appropriate chart type and complexity level to effectively convey the information [49].
+It is essential to adopt the appropriate chart type and complexity level to effectively convey the information (Evergreen, 2019).
 Previous works have focused mainly on charts with only a single data format and low information density, which are rarely encountered in practical settings such as academic writing.
-Noting that LMMs like GPT-4V have already demonstrated outstanding data visualization capabilities, our focus is on using charts that are actually employed in practice, such as those from research papers, and on selecting charts with varying levels of complexity when constructing our benchmark.
+Noting that LMMs like GPT-4o have already demonstrated outstanding data visualization capabilities, our focus is on using charts that are actually employed in practice, such as those from research papers, and on selecting charts with varying levels of complexity when constructing our benchmark.
 Reduction of Data Leakage.
@@
 To further illustrate the potential data leakage in matplotlib gallery, we select the code for $20$ charts from the matplotlib gallery.
-Specifically, we provide the first half of the code as a prefix for Llama3-8B [15] and let the model complete the remaining code.
+Specifically, we provide the first half of the code as a prefix for Llama3-8B (AI@Meta, 2024) and let the model complete the remaining code.
 Then we calculate the edit distance between the generated complete code block and the ground truth.
@@
 Users from diverse domains such as finance, healthcare, education, and engineering demonstrate unique needs and preferences for data visualization.
-These requirements extend beyond mere diversity; they demand the incorporation of charts capable of articulating complex and multi-dimensional data, and conforming to domain-specific aesthetic preferences[50, 49].
+These requirements extend beyond mere diversity; they demand the incorporation of charts capable of articulating complex and multi-dimensional data, and conforming to domain-specific aesthetic preferences(Qin et al., 2020; Evergreen, 2019).
 By aligning our dataset construction process with these real-world demands, we enable a more relevant and precise evaluation of LMMs.
@@
 This focused development will also lead to an enhanced user experience and increased user satisfaction, as the models become more proficient in meeting the sophisticated and varied needs of users.
-A.2 Data Annotation Pipeline
+B.2 Data Annotation Pipeline
+Figure 5: An illustration of the data annotation pipeline, which encompasses five key steps.
+We present the detailed description about the five-step data annotation pipeline in this section. Fig. 5 demonstrates an illustration of the data annotation pipeline.
 General Filtering.
 To obtain a high-quality dataset aligned with real-world use cases and to avoid data leakage, we initially scrape figures from source files of publications on arXiv [^3] https://arxiv.org/ that hold a CC BY 4.0 license and have a publication date after February 2024.
-We then extract figures in PDF format that contain at least one chart, yielding approximately $174,100$ figures across various domains (such as Physics, Mathematics, Computer Science, Quantitative Biology, Quantitative Finance, Statistics, Electrical Engineering and Systems Science, Economics).
-We further filter the figures based on criteria of how designers select inspiring visualization examples, including clarity and visual appeals[13, 14], color schemes[13], and uniqueness of the chart within its category, resulting in a refined collection of about $15,800$ charts.
-It is noteworthy that during this process, some charts, which do not entirely meet the clarity and aesthetic criteria, are retained in the corpus for their unique and rare qualities, pending further modifications.
+We then extract figures in PDF format, yielding approximately 174,100 figures across various domains (such as Physics, Mathematics, Computer Science, Quantitative Biology, Quantitative Finance, Statistics, Electrical Engineering and Systems Science, Economics).
+We filter these figures based on file format and generation method, retaining only Matplotlib-generated PDFs, indicating that these figures can be reproduced using Python.
+This process results in a refined collection of 15,800 figures.
 Diversity and Information Density Filtering.
-In this phase, we adopt a two-stage filtering process focused on the diversity of the charts and the effectiveness of information communication, primarily assessed based on the complexity of visual elements and the information density of each chart[13].
-Specifically, during the first round, we maintain a data pool, categorizing chart types and subcategories.
-Each chart is compared against the most similar one within the corresponding category in the data pool.
-If significant differences in visual complexity or information density are observed, the chart is added to the pool.
-Otherwise, we retain the one that better fulfills the criteria for diversity and effective information communication.
-After the first round of filtration, we observe a decreasing frequency of charts with increasing levels of difficulty among the remaining $15,800$ entries.
-Charts with simpler designs and lower information density often share greater similarity in their arrangement of visual elements, leading to a higher rate of duplicates being identified and removed.
-In contrast, more complex charts with higher information density are less likely to have comparable counterparts, thereby reducing duplication and enhancing the retention rate of these complex charts.
-This process effectively refines the complexity spectrum of our dataset, ensuring a balanced distribution.
-Subsequently, we conduct the second round of filtration.
-We engage five experts from different fields—computer science, visual communication design, and data visualization—to independently assess the data, creating five separate data pools.
-We then preserve the intersection of these pools and conduct a final, collective filtration from the combined set.
-The selection of which charts are to be included in the final pool is made through a voting process involving all five experts.
-Ultimately, this approach results in a curated collection of 279 charts.
+This stage involves a two-phase process conducted by five domain experts from information visualization, digital media, industrial design, visual communication design, and computer science.
+The information visualization expert has three years of research experience in scientific visualization and visual analytics.
+The digital media specialist focuses on human-computer interaction and multimedia design.
+The industrial design expert brings perspectives from user experience and product visualization.
+The visual communication design expert specializes in graphic design principles and information aesthetics.
+The computer science expert has extensive experience in data visualization programming and scientific computing.
+In the first phase, these experts conduct a 7-day manual review of the 15,800 figures, focusing on visual diversity and information communication effectiveness.
+They reference the Matplotlib gallery, gradually identifying and finalizing chart type while reviewing the figures, and build corresponding type pools.
+For each new figure, they assess its visual elements—such as layout, axes, line styles, marker styles, and colors—against existing figures in their corresponding type pool.
+If the figure as long as exhibits a distinctive difference in at least one of these aspects, it is retained; otherwise, it is excluded.
+This process results in 1,295 figures being selected for the second phase.
+In the second phase, these experts independently review the 1,295 figures and further select those figures they deem to exhibit significant distinctions and diversity.
+Figures selected unanimously by all experts are directly included, while the remaining figures are subjected to a majority voting system requiring at least 3/5 votes for inclusion.
+This rigorous process, which takes less than 3 days to complete, results in a final set of 279 figures.
 Manual Selection and Filling.
-In addition sourcing from arXiv, we curate chart figures from diverse platforms such as the matplotlib gallery, Stack Overflow, and plotting-related forums on Twitter and Reddit.
+In addition to sourcing from arXiv, we curate chart figures from diverse platforms such as the matplotlib gallery, Stack Overflow, and plotting-related forums on Twitter and Reddit.
 These charts are deliberately chosen for their distinctive styles, which are not present in our arXiv dataset.
-To prevent data leakage, we meticulously process the data and styling of these charts, substituting existing style schemes and data with those not present in our dataset, while preserving their visual appeal.
+Consequently, we obtain 600 prototype charts for ChartMimic.
+This stage took us less than a week to complete the data selection.
 Code and Instruction Writing.
-We propose to manually write codes and instructions for ChartMimic based on the collected $500$ prototype charts.
-To ensure the quality of the annotation results, we assemble a team of skilled Python users.
-These individuals are master students from computer science-related disciplines, with at least 4 years of experience using Python and extensive familiarity with matplotlib.
-Initially, the annotators are tasked with meticulously reproducing the $500$ prototype charts using Python and the matplotlib package, resulting in a set of $500$ (figure, code, instruction) triplets for the Direct Mimic task.
-Subsequently, to simulate the real-world scenario of Customized Mimic, annotators are instructed to modify the chart data in the Direct Mimic task by integrating new data from various domains.
-They are then required to modify the corresponding codes and instructions, leading to the $500$ (figure, code, instruction) triplets for the Customized Mimic task.
-Consequently, we obtain the ChartMimic benchmark, comprising $1,000$ high quality test examples.
-Appendix B Chart Taxonomy
+We propose to manually write codes and instructions for ChartMimic based on the collected 600 prototype charts.
+To ensure annotation quality, a team of skilled Python users—Python annotators—master’s students in computer science with 6+ years of Python and matplotlib experience—reproduce 600 prototype charts using Python 3.9.0 and matplotlib v3.8.4.
+Since the unannotated data in the figures cannot be fully restored, they can only be approximated when writing the code.
+This process generates 600 (figure, code, instruction) triplets for the Direct Mimic task and another 600 triplets for the Customized Mimic task by integrating data from various domains into the corresponding code and instructions, comprising 1,200 high-quality seed data.
+Data Augmentation.
+After developing the seed triplets, we proceed with a manual data augmentation process.
+Python Annotators modify various elements of each seed triplet, including data, color schemes, and mark styles, to create augmented triplets.
+For each seed triplet, we generate three additional augmented triplets.
+This process enhances our dataset, resulting in a total of $4,800$ triplets.
+The "Code and Instruction Writing" and "Data Augmentation" stages together take the data annotators approximately 1.5 months to complete.
+B.3 Complexity Levels
+Our complexity assessment framework is established upon two fundamental criteria, designed to comprehensively evaluate both the visual and technical aspects of chart generation:
+Visual Elements Complexity We systematically evaluate the sophistication of each chart through a comprehensive scoring mechanism applied to its constituent visual components.
+This encompasses the chart typology, data grouping structures, marker configurations, textual elements, chromatic schemes, compositional layouts, and coordinate systems.
+Each visual component is quantitatively assessed based on both its frequency and sophistication level (designated as 1-3 points for low/medium/high complexity, respectively).
+This multi-dimensional scoring approach ensures a thorough evaluation of the visual complexity inherent in each chart.
+Implementation Complexity We incorporate code complexity as a quantitative metric, measured primarily through code length and structural intricacy.
+This parameter effectively captures the technical sophistication required for accurate chart reproduction, including the complexity of data preprocessing, visualization logic, and stylistic customizations.
+The implementation complexity provides insights into the programming challenges associated with each chart type.
+Consequently, charts in ChartMimic are systematically categorized into three distinct complexity levels, each representing a specific combination of visual and implementation challenges:
+Easy: Fundamental chart configurations featuring minimal visual complexity and straightforward implementation requirements (e.g., in LABEL:fig:easy-case, monochromatic bar charts with sparse textual elements and simplified data representation).
+Medium: Charts exhibiting intermediate complexity in visual element composition or implementation requirements (e.g., in LABEL:fig:medium-case, dual-subplot bar charts incorporating grouped data structures, diverse chromatic schemes, and moderate textual annotations).
+Hard: Charts demonstrating sophisticated visual elements or advanced implementation (e.g., in LABEL:fig:hard-case, complex dual-subplot bar charts featuring divergent data patterns, extensive color schemes, comprehensive textual annotations, and substantial code complexity).
+B.4 Instruction Examples
+To illustrate our instruction format and task requirements, we present examples (bar_28 and CB_29) with their corresponding (figure, instruction, code) triplets for both Direct Mimic and Customized Mimic tasks.
+Fig.
+8 and Fig.
+9 show the Direct Mimic and Customized Mimic tasks for a bar chart (bar_28), respectively.
+The instruction provides guidance on creating the visualization.
+Similarly, Fig.
+10 and Fig.
+11 demonstrate the tasks for a more complex combination chart (CB_29).
+Figure 8: An Example of the Direct Mimic task (bar_28), showing the (figure, instruction, code) triplet. Additionally, we also display the ground truth figure rendered from the code for illustration.
+Figure 9: An Example of the Customized Mimic task (bar_28), showing the (figure, instruction, code) triplet.
+Additionally, we also display the ground truth figure rendered from the code for illustration.
+Figure 10: An Example of the Direct Mimic task (CB_29), showing the (figure, instruction, code) triplet. Additionally, we also display the ground truth figure rendered from the code for illustration.
+Figure 11: An Example of the Customized Mimic task (CB_29), showing the (figure, instruction, code) triplet.
+Additionally, we also display the ground truth figure rendered from the code for illustration.
+Appendix C Correlation Between Test Set and Testmini set
+Table 7: The ChartMimic benchmark with Direct Mimic task. We report the results of both Testmini and Test set.
+Model Params Test set Exec. Rate Low-Level High-Level Overall
+Text Layout Type Color Avg. GPT-4o
+Proprietary
+GPT-4o - Testmini 93.2 81.5 89.8 77.3 67.2 79.0 83.5 81.2
+Test 93.0 83.4 90.2 76.7 66.0 79.1 84.1 81.6
+Open-Weight
+InternVL2-2B 2.2B Testmini 52.5 23.6 35.8 16.0 15.4 22.7 24.2 23.5
+Test 51.8 23.3 34.8 15.9 15.9 22.5 25.6 24.0
+InternVL2-8B 8.1B Testmini 61.8 31.5 51.1 28.6 26.2 34.4 38.9 36.6
+Test 61.6 33.4 47.4 25.7 24.2 32.7 39.9 36.3
+InternVL2-26B 26.0B Testmini 69.3 39.2 58.7 35.9 31.8 41.4 47.4 44.4
+Test 69.9 41.7 58.0 35.6 31.0 41.6 48.1 44.9
+InternVL2-Llama3-76B 76.0B Testmini 83.2 54.1 74.5 49.2 41.5 54.8 62.2 58.5
+Test 83.3 55.6 73.5 50.4 40.8 55.1 62.7 58.9
+Table 8: The ChartMimic benchmark with Customized Mimic task. We report the results of both Testmini and Test set.
+Model Params Test set Exec. Rate Low-Level High-Level Overall
+Text Layout Type Color Avg. GPT-4o
+Proprietary
+GPT-4o - Testmini 96.5 88.5 92.9 79.2 67.6 82.1 84.3 83.2
+Test 96.2 87.2 91.7 80.1 66.4 81.4 84.8 83.1
+Open-Weight
+InternVL2-2B 2.2B Testmini 49.3 22.2 35.4 20.0 18.1 23.9 27.8 25.9
+Test 49.6 22.4 33.9 19.2 19.6 23.8 28.4 26.1
+InternVL2-8B 8.1B Testmini 73.0 43.1 54.4 39.9 35.4 43.2 48.9 46.1
+Test 73.5 43.7 54.1 41.1 34.1 43.3 49.8 46.5
+InternVL2-26B 26.0B Testmini 73.7 43.9 62.3 43.5 34.3 46.0 51.1 48.6
+Test 74.7 44.7 66.3 46.8 35.1 48.2 50.8 49.5
+InternVL2-Llama3-76B 76.0B Testmini 89.8 57.8 79.0 63.5 50.5 62.7 66.7 64.7
+Test 88.1 57.9 79.6 65.6 51.7 63.7 68.2 66.0
+Tab.
+7 and 8 reports the performance of GPT-4o and InternVL2 series LMMs (2B, 8B, 26B, Llama3-76B) on Direct Mimic and Customized Mimic task.
+The minor differences between scores on the test subset and the testmini subset suggest that testmini effectively mirrors the test subset, serving as a valuable evaluation subset for model development, especially for those who have limited computing resources.
+Appendix D Chart Taxonomy
 This section presents the chart taxonomy in ChartMimic. It encompasses a structure of 22 categories according to chart type characteristics and data composition. The categories comprise of:
@@
 Derived from the categories above and using diverse layouts, Multidiff charts offer numerous configurations, as shown in Fig.
-24.
+31.
 Figure 31: Examples of Multidiff charts.
@@
 Additional examples are shown in Fig.
-25.
+32.
 Figure 32: Examples of Combination charts.
@@
 Figure 33: Examples of HR charts.
-Appendix C Details of Evaluation Metrics
-In this section, we present the details of evaluation metrics, including GPT-4V score for high-level metrics and Text Score, Layout Score, Type Score and Color Score for low-level metrics.
+Appendix E Details of Evaluation Metrics
+In this section, we present the details of evaluation metrics, including GPT-4o score for high-level metrics and Text Score, Layout Score, Type Score and Color Score for low-level metrics.
 For more information, please refer to our source code.
-C.1 GPT-4V Score
-We employ GPT-4V [2] to assess the extend to which the generated figure corresponds to the ground-truth figure.
+E.1 GPT-4o Score
+We employ GPT-4o (OpenAI, 2024) to assess the extend to which the generated figure corresponds to the ground-truth figure.
 The specific content of the prompt is presented in Fig.
-27.
-Specifically, we input both the generated and the ground-truth figures into the GPT-4V simultaneously.
-Then, GPT-4V is instructed to evaluate the similarity between the two figures, taking into account six dimensions: text, layout, type, data, style, and clarity.
-Subsequently, GPT-4V outputs a score ranging from 0 to 100 to represent the degree of similarity between the figures.
-Prompt for GPT-4V Score You are an excellent judge at evaluating visualization chart plots.
+34.
+Specifically, we input both the generated and the ground-truth figures into the GPT-4o simultaneously.
+Then, GPT-4o is instructed to evaluate the similarity between the two figures, taking into account six dimensions: text, layout, type, data, style, and clarity.
+Subsequently, GPT-4o outputs a score ranging from 0 to 100 to represent the degree of similarity between the figures.
+Prompt for GPT-4o Score You are an excellent judge at evaluating visualization chart plots.
 The first image (reference image) is created using ground truth matplotlib code, and the second image (AI-generated image) is created using matplotlib code generated by an AI assistant.
@@
 Use the following format for your response: Comments: - Chart Types: \${your comment and subscore} - Layout: \${your comment and subscore} - Text Content: \${your comment and subscore} - Data: \${your comment and subscore} - Style: \${your comment and subscore} - Clarity: \${your comment and subscore} Score: \${your final score out of 100} Please use the above format to ensure the evaluation is clear and comprehensive.
-Figure 27: Prompt for GPT-4V Score.
-C.2 Text Score
+Figure 34: Prompt for GPT-4o Score.
+Stability of Evaluation with GPT-4o Score.
+We conduct the high-level evaluation for GPT-4o on the Direct Mimic task for 5 times to assess the stability of GPT-4o Score.
+The result indicate a mean GPT-4o Score of $83.4$ with a standard deviation of 0.08, demonstrating the stability of GPT-4o Score.
+Cost of Evaluation with GPT-4o Score.
+A single-round evaluation with GPT-4o Score on the Direct Mimic task approximately costs \$5.25.
+Utilizing OpenAI’s batch services further reduces this cost to \$2.63 per round.
+Longevity of Evaluation with GPT-4o Score.
+The GPT-4o Score metric is designed with long-term viability in mind.
+Our approach provides a meta-evaluation framework that can adapt to evolving language models.
+While currently leveraging GPT-4o, the method is model-agnostic and can be implemented with other advanced LLMs such as Claude or Gemini.
+This adaptability ensures that as more capable models emerge, the evaluation process can seamlessly transition to utilize these superior systems.
+Importantly, this approach has demonstrated increasing correlation with human assessments as model capabilities improve, as evidenced by studies like AlpacaEval (Li et al., 2023) and MT-Bench (Zheng et al., 2023).
+Figure 35: Example of GPT-4o’s scoring results for a Direct Mimic example (bar_28).
+Figure 36: Example of GPT-4o’s scoring results for a Customized Mimic example (bar_28).
+Figure 37: Example of GPT-4o’s scoring results for a Direct Mimic example (CB_29).
+Figure 38: Example of GPT-4o’s scoring results for a Customized Mimic example (CB_29).
+GPT-4o Score Examples.
+To demonstrate our evaluation framework, we present GPT-4o’s scoring examples for both Direct Mimic and Customized Mimic tasks.
+Figs.
+35, 36, 37 and 38 show the evaluation results for two representative charts of different complexity levels (bar_28 and CB_29).
+These examples demonstrate how GPT-4o systematically evaluates various aspects of chart reproduction across different chart complexities and task types.
+Correlation Coefficient Comparison with CLIP Score.
+CLIP Score (Radford et al., 2021) is widely used for assessing figure similarity.
+However, our preliminary experiments indicate that it struggles to differentiate variations in types and other critical elements in charts, resulting in a low correlation coefficient of 0.53 with human evaluation.
+In contrast, as stated in Sec.
+4.3, GPT-4o Score achieves a correlation coefficient of 0.70 with human evaluation.
+Therefore, we adopt GPT-4o Score as the high-level evaluation metric.
+E.2 Text Score
 Listing 1: An exemplary Python code for logging text information.
@@
 Subsequently, we employ this approach to extract the text information from both the generated code and the ground-truth code.
-The F1 score, computed between these two groups of texts, is denoted as Text Score.
-C.3 Layout Score
+Based on these two groups of texts, precision is defined as the ratio of the number of correctly captured ground-truth text to the total number of text in the generated figure.
+Recall is defined as the ratio of the number of correctly captured ground-truth text to the total number of text in the ground-truth figure.
+The F1-score, calculated using precision and recall, serves as Text Score.
+E.3 Layout Score
 Listing 2: An exemplary Python code for logging layout information.
@@
 An exemplary Python code to accomplish this process is provided in Listing LABEL:lst:layout, where the position information of each axis is recorded.
-Subsequently, we gather the layout information from both the ground-truth code and the generated code, and compute the F1 score between the them.
-This metric is referred as the Layout Score.
-C.4 Type Score
+Subsequently, we gather the layout information from both the ground-truth code and the generated code.
+Based on these two groups of layout information, precision is defined as the ratio of the number of correctly identified ground-truth layouts to the total number of layouts in the generated figure.
+Recall is defined as the ratio of the number of correctly identified ground-truth layouts to the total number of layouts in the ground-truth figure.
+The F1-score, calculated using precision and recall, serves as Layout Score.
+E.4 Type Score
 Listing 3: An exemplary Python code for logging type information.
@@
 Listing LABEL:lst:charttype presents an exemplary Python code that demonstrates this approach, where we introduce a logger to the “bar()” function.
-Subsequently, we gather the chart types from both the generated and ground-truth code, and compute the F1 score between them.
-We refer to this metric as Type Score.
-C.5 Color Score
+Subsequently, we gather the chart types from both the generated and ground-truth code.
+Based on these two groups of chart types, precision is defined as the ratio of the number of correctly identified ground-truth chart types to the total number of chart types in the generated figure.
+Recall is defined as the ratio of the number of correctly identified ground-truth chart types to the total number of chart types in the ground-truth figure.
+The F1-score, calculated using precision and recall, serves as Type Score.
+E.5 Color Score
 Listing 4: An exemplary Python code for logging color information.
@@
 An example Python code demonstrating this approach is provided in Listing LABEL:lst:color.
-Once we obtain the color attributes from both the ground-truth code and the generated code, we compare them using the F1 score metric.
+Subsequently, we gather the color information from both the ground-truth code and the generated code and calculate the similarity between them.
 It is noteworthy that through preliminary experiments, we find that using exact color matching resulted in very low color similarity, as even slight variations would result in a similarity score of zero.
-To address this issue, we employ the CIEDE2000 color difference formula[51], which converts the matching value between two colors from a discrete [0,1] scale to a continuous range between 0 and 1.
-Finally, we calculate the maximum color similarity between the sets of colors in the ground-truth code and the generated code to compute F1 score.
-We refer to the F1 score as Color Score.
-Appendix D Generation Configurations and Prompting Methods
-D.1 Generation Configurations
-Following previous setups [52, 53], for open-weight models, we set the temperature $\tau=0.1$ to achieve optimal results, while for proprietary models, we set the temperature $\tau=0$ for greedy decoding.
+To address this issue, we employ the CIEDE2000 color difference formula (Luo et al., 2001), which converts the matching value between two colors from a discrete $[0,1]$ scale to a continuous range between 00 and $1$.
+Finally, we calculate the maximum color similarity between the sets of colors in the ground-truth code and the generated code.
+Precision is defined as the ratio of maximum color similarity to the total number of color in the generated figure.
+Recall is defined as the ratio of maximum color similarity to the total number of color in the ground-truth figure.
+The F1-score, calculated using precision and recall, serves as Color Score.
+Appendix F Model Configurations and Prompting Methods
+F.1 Generation Configurations
+Following previous setups (Wang et al., 2023b; Shi et al., 2024), for open-weight models, we set the temperature $\tau=0.1$ to achieve optimal results, while for proprietary models, we set the temperature $\tau=0$ for greedy decoding.
 For all models, we set the maximum generation length to 4096.
 Additionally, we use BF16 for model inference for open-weight models.
-All models are inferred on a single A100 80G GPU.
-D.2 Prompts
-We provide prompts for Direct, HintEnhanced, SelfRevision and Scaffold Prompting in Figs.
-28, 29, 30, 31 and 32.
+All models are inferred on A100 80G GPU.
+F.2 Prompts
+We provide prompts for Direct, HintEnhanced, SelfReflection and Scaffold Prompting in Figs.
+39, 40, 41, 43 and 42.
 for Direct Prompting, we meticulously design a separate prompt for open-source models to achieve optimal results, as shown in Fig.
-28.
+39.
 while the prompt for proprietary Models is shown in Fig.
-29.
+40.
 Prompt for Direct Prompting (Open-Weight Models) You are an expert Python developer who specializes in writing matplotlib code based on a given picture.
@@
 I need your help to generate the Python code that can reproduce the picture based on the picture I provide.
-Please note that it is necessary to use figsize=({width} (width), {height} (height)) to set the image size to match the original size.
+Please note that it is necessary to use figsize=({width}, {height}) to set the image size to match the original size.
 Additionally, I will not provide you with the actual data in the image, so you have to extract the actual data by yourself and based on the extracted data to reproduce the image.
@@
 I need your help to generate the Python code that can reproduce the picture based on the picture I provide.
-Note that it is necessary to use figsize=({width} (width), {height} (height)) to set the image size to match the original size.
+Note that it is necessary to use figsize=({width}, {height}) to set the image size to match the original size.
 Now, please give me the matplotlib code that reproduces the picture below.
@@
 Now, given the picture below, please first output your comprehensive caption and then use the caption to assist yourself to generate matplotlib code that reproduces the picture.
-Note that it is necessary to use figsize=({width} (width), {height} (height)) to set the image size to match the original size.
+Note that it is necessary to use figsize=({width}, {height}) to set the image size to match the original size.
 Figure 41: Prompt for HintEnhanced Prompting. {text} in blue font represents placeholders, which varies according to different test examples.
-Prompt for SelfRevision Prompting You are an expert Python developer who specializes in writing matplotlib code based on a given picture.
-I have a code for implementing the reference picture as follows: Now, I have the Python matplotlib code for implementing the reference picture as follows: ```python {python_function} ``` The rendered picture of the code is: Now, please compare whether the renderer picture is the same as the reference picture.
+Prompt for Scaffold Prompting You are an expert Python developer who specializes in writing matplotlib code based on a given picture.
+I found a very nice picture in a STEM paper, but there is no corresponding source code available.
+I need your help to generate the Python code that can reproduce the picture based on the picture I provide.
+I will provide you with two images.
+The first image is the original picture.
+The second image is the picture overlaid with a dot matrix of a shape of {dot_matrix_height} * {dot_matrix_width} to help you with your task, and each dot is labeled with two-dimensional coordinates (x,y).
+Within each column, the x-coordinate increases from top to bottom, and within each row, the y-coordinate increases from left to right.
+Please first use this dot matrix as reference anchors to generate the description of the picture (e.g., between dot A and dot B is something) and generate matplotlib code that reproduces the picture.
+Note that it is necessary to use figsize=({width}, {height}) to set the image size to match the original size.
+Figure 42: Prompt for Scaffold Prompting. {text} in blue font represents placeholders, which varies according to different test examples.
+Prompt for SelfReflection Prompting You are an expert Python developer who specializes in writing matplotlib code based on a given picture.
+I have a code for implementing the reference picture as follows: Now, I have the Python matplotlib code for implementing the reference picture as follows: ```python {python_code} ``` The rendered picture of the code is: Now, please compare whether the renderer picture is the same as the reference picture.
 The difference may cover, but not be limited to, the following aspects: 1.
@@
 Finally, output the revised code.
-Note that it is necessary to use figsize=({width} (width), {height} (height)) to set the image size to match the original size.
-Figure 31: Prompt for SelfRevision Prompting. {text} in blue font represents placeholders, which varies according to different test examples.
-Prompt for Scaffold Prompting You are an expert Python developer who specializes in writing matplotlib code based on a given picture.
-I found a very nice picture in a STEM paper, but there is no corresponding source code available.
-I need your help to generate the Python code that can reproduce the picture based on the picture I provide.
-I will provide you with two images.
-The first image is the original picture.
-The second image is the picture overlaid with a dot matrix of a shape of {dot_matrix_height} * {dot_matrix_width} to help you with your task, and each dot is labeled with two-dimensional coordinates (x,y).
-Within each column, the x-coordinate increases from top to bottom, and within each row, the y-coordinate increases from left to right.
-Please first use this dot matrix as reference anchors to generate the description of the picture (e.g., between dot A and dot B is something) and generate matplotlib code that reproduces the picture.
-Note that it is necessary to use figsize=({width} (width), {height} (height)) to set the image size to match the original size.
-Figure 32: Prompt for Scaffold Prompting. {text} in blue font represents placeholders, which varies according to different test examples.
-D.3 Error Cases of Different Prompting Methods
-We provide error cases of HintEnhanced, SelfRevision and Scaffold Prompting in Figs. 33, 34 and 35. For an analysis of error cases related to Direct Prompting, please refer to Appendix F.
-Figure 33: An error case of HintEnhanced Prompting. The text highlighted in blue is GPT-4V’s incorrect understanding regarding provided figures. The remaining part of response is omitted with “……”.
-Figure 34: An error case of SelfRevision Prompting. The text highlighted in blue is GPT-4V’s incorrect understanding regarding provided figures. The remaining part of response is omitted with “……”.
-Figure 35: An error case of Scaffold Prompting. The text highlighted in blue is GPT-4V’s incorrect understanding regarding provided figures. The remaining part of response is omitted with “……”.
-Appendix E Correlation with Human Evaluation
+Note that it is necessary to use figsize=({width}, {height}) to set the image size to match the original size.
+Figure 43: Prompt for SelfReflection Prompting. {text} in blue font represents placeholders, which varies according to different test examples.
+F.3 Cases of Different Prompting Methods
+We provide cases of HintEnhanced, SelfReflection and Scaffold Prompting in Figs. 44, 45, 46 and 47. For an analysis of error cases related to Direct Prompting, please refer to Appendix I.
+Figure 44: An error case of HintEnhanced Prompting. The text highlighted in red is LMM’s incorrect understanding regarding provided figures. The remaining part of response is omitted with “……”.
+Figure 45: An error case of SelfReflection Prompting. The text highlighted in red is LMM’s incorrect understanding regarding provided figures. The remaining part of response is omitted with “……”.
+Figure 46: An good case of SelfReflection Prompting. The text highlighted in blue is LMM’s correct reflection regarding provided figures. The remaining part of response is omitted with “……”.
+Figure 47: An error case of Scaffold Prompting. The text highlighted in red is LMM’s incorrect understanding regarding provided figures. The remaining part of response is omitted with “……”.
+F.4 Details of Models
+We list the evaluated models in Tab. 9.
+Model Params Language Model Vision Model Model Code/API
+GPT-4o - - - gpt-4o-2024-05-13
+Claude-3-opus - - - claude-3-opus-20240229
+GeminiProVision - - - gemini-pro-vision
+InternVL2-Llama3-76B 76.0B Llama-3-70B-Instruct InternViT-6B OpenGVLab/InternVL2-Llama3-76B
+LLaVA-Next-Yi-34B 34.8B Yi-34B CLIP ViT-L/14 llava-hf/llava-v1.6-34b-hf
+InternVL2-26B 26.0B InternLM2-20B InternViT-6B OpenGVLab/InternVL2-26B
+Cogvlm2-llama3-chat-19B 19.2B Llama-3-8B-Instruct EVA2-CLIP-E THUDM/cogvlm2-llama3-chat-19B
+Phi-3-Vision-128K-Instruct 4.2B Phi-3 CLIP ViT-L/14 microsoft/Phi-3-vision-128k-instruct
+IDEFICS2-8B 7.6B Mistral-7B SigLip-400M HuggingFaceM4/idefics2-8b
+LLaVA-Next-Mistral-7B 7.6B Mistral-7B CLIP ViT-L/14 llava-hf/llava-v1.6-mistral-7b-hf
+DeepSeek-VL-7B 7.3B DeekSeek-7B SAM-B & SigLIP-L deepseek-ai/deepseek-vl-7b-chat
+MiniCPM-Llama3-V2.5 8.4B Llama3-8B-Instruct SigLip-400M openbmb/MiniCPM-Llama3-V-2_5
+Qwen2-VL-7B 8.2B Qwen2-7B ViT-600M Qwen/Qwen2-VL-7B-Instruct
+InternVL2-8B 8.1B InternLM2.5-7B InternViT-300M OpenGVLab/InternVL2-8B
+InternVL2-4B 4.2B Phi-3 InternViT-300M OpenGVLab/InternVL2-4B
+Qwen2-VL-2B 2.6B Qwen2-1.5B ViT-600M Qwen/Qwen2-VL-2B-Instruct
+InternVL2-2B 2.2B InternLM2-1.8B InternViT-300M OpenGVLab/InternVL2-2B
+Table 9: Model code/API of the evaluated models.
+Appendix G Correlation with Human Evaluation
 Figure 48: A screenshot of the human evaluation questionnaire.
@@
 Specifically, a selection of $300$ test examples from the Direct Mimic task is utilized.
-Subsequently, we gather the results generated by GPT-4V using four different prompting methods (Sec.
+Subsequently, we gather the results generated by GPT-4o using four different prompting methods (Sec.
 4.2) on these test examples.
@@
 A screenshot of the final questionnaire is available in Fig.
-36.
+48.
 Upon collecting the human evaluation results, we calculate the Pearson correlation coefficient to ascertain the correlation between our automatic metrics and human evaluation.
-Appendix F Cases of Error Analysis
+Appendix H Correlation with Chart Understanding and Code Generation
+Figure 49: Performance correlation of ChartMimic with benchmarks assessing chart understanding and code generation capabilities.
+In an effort to shed light on the factors enhancing performance on ChartMimic, we explore the performance correlation between ChartMimic and existing benchmarks that evaluate chart understanding and code generation capabilities.
+Specifically, we select ChartQA (Masry et al., 2022) as the benchmark for chart understanding and HumanEval (Chen et al., 2021) as the benchmark for code generation.
+We then calculate two Pearson correlation coefficients to quantify these relationships: The correlation coefficient between the performance of LMMs in the Direct Mimic task and their performance in the ChartQA task, denoted as $r_{chart}$, and the correlation coefficient between the performance of LMMs in the Direct Mimic task and the performance of their corresponding LLMs in the HumanEval task, denoted as $r_{code}$
+As depicted in Fig.
+49, the calculated values are $r_{chart}$ = 0.6877 and $r_{code}$ = 0.8714.
+These results indicate that both chart understanding and code generation abilities influence the performance in the Direct Mimic task, with code generation having a more significant impact.
+This implies that for future model development aimed at enhancing multimodal code generation capabilities, it is crucial to focus on the foundational code generation abilities of LMMs.
+Appendix I Cases of Error Analysis
 We provide cases of text-related, type-related and color-related errors in LABEL:fig:error_1, LABEL:fig:error_2, LABEL:fig:error_3, LABEL:fig:error_4, LABEL:fig:error_5 and LABEL:fig:error_6.
@@
 4.4.
-Appendix G Ethics and Societal Impact
-We envision ChartMimic as a comprehensive suite of benchmarks designed to guide researchers in assessing the capabilities of their LMMs.
-By providing an extensive evaluation framework, ChartMimic aims to foster the development of foundation models for the community, offering insights into various aspects of LMMs, such as visual understanding, code generation, and cross-modal reasoning.
-When constructing prototype charts, we use figures from papers on arXiv, which are licensed under CC BY 4.0.
-During the manual annotation of the code, we implement strict ethical controls to ensure that the information conveyed in the code does not contain any biased content and does not infringe on any private data.
+Appendix J Error Analysis of Open-Weight Models
+Figure 51: Error analysis of InternVL2-Llama3-76B across four error types on the Direct Mimic task. The number in brackets indicates the count of error case.
+Figure 52: Error analysis of InternVL2-26B across four error types on the Direct Mimic task. The number in brackets indicates the count of error case.
+Figure 53: Error analysis of DeepSeek-VL-7B across four error types on the Direct Mimic task. The number in brackets indicates the count of error case.
+We conduct detailed error analysis on representative open-weight models of different scales, sampling 100 cases for each model.
+For large-scale models such as InternVL2-Llama3-76B and InternVL2-26B (Fig.
+51Fig.
+52), the error patterns align closely with those observed in GPT-4o.
+Specifically, Code-related Errors predominantly arise from dimensional inconsistencies, while Text-related Errors primarily manifest as missing elements.
+Type-related Errors mainly comprise missing components cases, and Color recognition demonstrates the capability to identify similar, if not identical, chromatic properties.
+In contrast, smaller-scale models like DeepSeek-VL-7B (Fig.
+53) exhibit markedly different error distributions.
+Code-related issues show a significant increase in Access and Parameter Errors (approximately 53.7%), while Text-related errors demonstrate a higher prevalence of missing elements (approximately 66.7%).
+Type-related cases reveal a substantial increase in missing errors (approximately 78.4%).
+Moreover, the degradation in color recognition is more severe, with the model failing to identify even similar colors.
+These findings indicate that while larger models encounter challenges similar to GPT-4o, smaller models exhibit more fundamental difficulties in both code generation and visual comprehension.
+Appendix K Ethics, Societal Impact and Scalability of ChartMimic
+ChartMimic is designed as a comprehensive evaluation framework for advancing LMMs, with a strong emphasis on ethical considerations, societal impact and scalability.
+Its primary objective is to provide researchers with a rigorous, ethically sound tool for assessing LMMs’ capabilities across critical domains, including visual understanding, code generation and cross-modal reasoning.
+Ethical Considerations.
+The ethical integrity of ChartMimic is foundational to its design and implementation.
+Our dataset is derived from scientific domains, specifically using figures from arXiv papers distributed under the CC BY 4.0 license.
+This focus on scientific content significantly mitigates potential ethical concerns related to bias, representation, or sensitive personal information that often arise in datasets derived from social media or general web content.
+The scientific nature of our data ensures: Minimal risk of perpetuating societal biases or stereotypes; Absence of personally identifiable information; Content that is generally neutral and objective, focusing on factual scientific representations.
+Our manual annotation process for code generation incorporates stringent ethical controls.
+Annotators are trained to ensure that the information conveyed in the code does not contain any biased content and strictly adheres to the factual representation of the scientific charts.
+This process further reinforces the ethical robustness of our dataset.
+Societal Impact.
+By providing a benchmark rooted in scientific data, ChartMimic contributes to the development of LMMs that can better understand and interact with complex, data-driven visualizations.
+This capability has far-reaching positive implications for scientific communication, data analysis, and knowledge dissemination across various fields.
+Scalability.
+The architecture of ChartMimic is inherently extensible, featuring a modular codebase that facilitates the seamless integration of additional chart types and evaluation metrics.
+Our data collection methodology leverages the continuous update cycle of arXiv, enabling sustainable expansion of the dataset while maintaining its focus on ethically sound, scientific content.
+The evaluation framework also demonstrates notable scalability: The concept of low-level metric can be readily adapted to other programming environments such as JavaScript and R.
+The high-level metric incorporates a meta-evaluation approach using LMMs, which allows for sustainable alignment with human preferences as stronger models emerge.